### v2

Abstract
This paper introduces innovative enhancements to the K-means and K-nearest neighbors (KNN) algorithms based on the concept of Gini prametric spaces. Unlike traditional distance metrics, Gini prametrics incorporate both value-based and rank-based measures, offering robustness to noise and outliers. The paper’s main contributions include (1) proposing a Gini prametric that captures rank information alongside value distances, (2) presenting a Gini K-means algorithm that is proven to converge and demonstrates resilience to noisy data, and (3) introducing a Gini KNN method that rivals state-of-the-art approaches like Hassanat’s distance in noisy environments. Experimental evaluations on 14 datasets from the UCI repository reveal the superior performance and efficiency of Gini-based algorithms in clustering and classification tasks. This work opens new avenues for leveraging rank-based prametrics in machine learning and statistical analysis.
1 Introduction
Machine learning, data analysis, and pattern recognition are mainly divided into two fields corresponding to two different tasks. The former is employed for supervised classification (with labeled data), while the latter is used for unsupervised classification or clustering (with unlabeled data). In each field, numerous algorithms are available based on different metrics, however some distance metrics are used in both contexts, classification and clustering. For instance, K-nearest neighbors or KNN Fix & Hodges (1951); Cover & Hart (1967) and K-means algorithms Steinhaus (1957); Lloyd (1957; 1982); Ball & Hall (1965); MacQueen (1967) were, at the moment of their creation, based on the Euclidean distance, before moving to a more flexible metric i.e. the Minkowski -distance (Oti et al., 2021).
On the one hand, K-means involves assigning points to clusters by iteratively measuring the distance between the points and the centroids of the clusters, with centroids being updated at each step of the algorithm until convergence. Jain (2008; 2010) shows that more than one thousand algorithms are available for K-means clustering. The developments of K-means are either concerned with initialization, convergence and partitioning the data (see for instance Pelleg & Moore (1999); Steinbach et al. (2000)) or the distance metric to be used to measure the proximity or the similarity between points and centroids, see for example the Kernel K-means to incorporate non-linearity in distances Schölkopf et al. (1998), the Fuzzy C-means Eschrich et al. (2003), the K-menoids based on the median Kaufman & Rousseeuw (2005), the X-means with Akaike or Bayesian information criteria Pelleg & Moore (2000), and recently Fair K-means Bateni et al. (2024) including fairness in distance functions.
On the other hand, KNN allows assigning points of the testing dataset to labels by measuring the distance between these test points and the K labeled points in the training dataset that are closest. The algorithm is therefore dependent on the number of neighbors K, but also, and more importantly, dependent on the choice of the distance Syriopoulos et al. (2023). Alfeilat et al. (2019) show that non-convex distances are well suited for managing noise in data whereas Vincent & Bengio (2001) show that convex distances defined on non-linear manifolds may be a good strategy to refine the classification. As in the K-means, weighting the points differently Gini (1912) or by means of fuzzification may improve the accuracy of the algorithm Derrac et al. (2014). However, as mentioned in Alfeilat et al. (2019), comparing 44 distances brings out 8 distances that outperform the others on many datasets (see Section 2 for details).
Prametric (or premetric) spaces have been introduced by Arkhangelskii & Pontryagin (1990). From our knowledge, apart from the use of non-convex distances, prametrics have not been employed in classification. A prametric space is a topological space that is more general than metric spaces, requiring neither symmetry nor indistinguishability, nor the validity of triangular inequality. It is provided with a function that is more general than a “distance” in the usual sense, since must satisfy non-negativity and . is then said to be a “prametric”, and since prametrics are not metrics (they usually do not satisfy the triangle inequality), they do not induced metric spaces as usual in topology. Indeed, the topology of a space induced by a prametric is necessarily sequential.
The Gini index is often used in multivariate statistics because it is a rank-dependant statistics robust to noise and outliers. This robustness holds true for many tasks such as regressions Yitzhaki & Schechtman (2004); Yitzhaki & Lambert (2013) or compression Charpentier et al. (2021). We shown in this paper that a prametric Gini can be defined with an hyper-parameter and employed for supervised classification and clustering to deal with noisy data. The main contributions of this paper are threefold:
1. To propose a Gini prametric based both on rank features and values of the points to capture two kinds of information, in particular the rank vectors for robustness.
2. To show that the proposed Gini K-means algorithm is convergent and robust to noise.
3. To show that Gini KNN may compete with the KNN based on the Hassanat distance, which is considered to be the most robust to noise.
The paper is structured as follows: Section 2 introduces existing works related to distance metrics for K-means and KNN algorithms. Section 3 presents the new Gini prametric proposed and investigated in our work. Section 4 introduces Gini K-means and Gini KNN. Section 5 is dedicated to the evaluation protocol and outlines the performance of the Gini prametric compared to existing approaches. Section 6 discusses our findings before mentioning perspectives.
2 State of the art
Metric distances play a central role in both supervised and unsupervised learning methods. The choice of the distance is challenging in practice, because of their variety and also because of the difficulty of determining in advance which distance is the most appropriate for a given dataset. However, Alfeilat et al. (2019) demonstrate that two distances outperform other ones (among a set of 44 distances) when dealing with noisy datasets and the K-nearest neighbor (KNN) classifier: the Hassanat distance and distances based on the L1 distance (Manhattan or city block distance), such as the Lorentzian and the Canberra distances. Other studies on K-means show that robust clustering may be performed thanks to weighted Manhattan distances An & Jiang (2024), quantum Manhattan distances Wu et al. (2022).
A distance satisfies the following properties: non-negativity, symmetry, nullity if and only if the points are equal, and finally the triangle inequality. Apart from Elkan (2003) who shows that the triangle inequality may serve to accelerate the K-means algorithm, there is no justification for all four of these properties to be satisfied at the same time. It might be possible to resort to pseudo-distances or prametrics to look for more flexible functional forms. From the review made by Alfeilat et al. (2019), eight families of distances can be identified in the literature. Let be two points, we provide in Table 1 the eight families with one member of each family (or two) that receives attention from the literature.
| Minkowski (F1) | |
| L distance | |
| Manhattan | |
| L1 distances (F2) | |
| Lorentzian | |
| Canberra | |
| Inner product (F3) | |
| Cosine distance | |
| Squared Chord (F4) | |
| Hellinger | |
| Squared L2 (F5) | |
| Squared Chi squared | |
| Shannon entropy (F6) | |
| Jensen-Shannon | |
| Vicissitude (F7) | |
| Vicis Symmetric | |
| Others (F8) | |
| Hassanat | if : |
| else: | |
In particular, Alfeilat et al. (2019) show that Hassanat’s distance Hassanat (2014) outperforms 43 other distances of the eight families based on the average precision and recall over many datasets of the UCI machine learning repository. They perform the same experiment with noise (from 10% to 90% of noise in the data with increment of 10%) and find that the same ranking is obtained in terms of averaged recalls and precisions: 1) Hassanat, 2) Lorentzian (F2), 3) Canberra (F2). The F1 and F3 distance families are less efficient, particularly the Euclidean and Cosine distances, as well as the Pearson distance (in F8) based on Pearson’s correlation coefficient.
3 Gini prametric spaces
3.1 Variance, Covariance, Norm and Distance
Given a vector , its L2 norm is simply , induced by the natural inner product in , . And the associated L2 distance is , but if neither first order homogeneity nor the triangle equality are need, we could simply consider the prametric , i.e., . In statistics, the norm is associated with the “standard deviation”, and the inner product is the standard Pearson’s linear correlation, for centered vectors. For a centered random vector , one can define its Pearson-covariance matrix as .
Nevertheless, it is well established that those measures are sensitive to outliers. A classical strategy is then to move from the L2 norm () to the L1 norm (). E.g., instead of using the average (solution of ), one could consider the median (solution of ). And the L1 norm is related to ranks, since, if and denotes a real-valued random variable with c.d.f. , while . Then
if is absolutely continuous, while, if ,
when the empirical e.c.d.f. is based on ranks (if is ordered, where ), where denote the rank of among . The use of ranks has been intensively considered in statistics to avoid the influence of outliers, as discussed in Kendall (1948); Hollander (1973); Huber (1981). Hence, Spearman (1904) suggested to use Spearman’s covariance, where ranks are considered, instead of numerical values.
3.2 Gini Index and Gini Covariance
The popular Gini index is defined as a normalized version of “Gini Mean Difference”, Gini (1912), defined as the expected absolute difference between two randomly drawn observations from the same population. The later can be rewritten with the covariance operator,
And Schechtman & Yitzhaki (1987) introduced a Gini Covariance operator, defined, for a pair of variables, as
where that is centered, and therefore
This dispersion measure, , involves the usual covariance of one of the variables with the rank of the other one, as a compromise between Pearson covariance and the Spearman version, based on pure ranks, . Its empirical version is a -statistic. If it is (unfortunately) asymmetric, this function is homogeneous of order 1, in the sense that, for ,
And more generally, given a random vector in , its Gini-covariance matrix is then
which is a positive matrix (see Dang et al. (2019); Charpentier et al. (2021)). Thus, in dimension 2, we can define the Gini-covariance between vectors in as,
We can now use those dissimilarity measures on variables to define a prametric distance on the set of individuals.
3.3 Gini Prametric
Let some data with points and features. The (cumulative) rank vector of feature assigns to each element of vector (column) its position (in ascending order of ). Inspired by the prametric defined as the squared Euclidean L2 distance,
we can define Gini prametric as follows:
Definition 3.1.
The Gini prametric (or Gini distance) is defined for all such that and are two given rows of (denoted “”):
where denotes the ascending rank of element in .
In practice, the average rank method is usually employed to deal with ties between many points, in this case the arithmetic mean of their ranks is computed. This avoids points with same values to be weighted by different ranks (this would imply for instance a bias in the Gini index, see Yitzhaki & Schechtman (2013)).
This Gini prametric is the empirical version of the following function defined for all , seen as realizations of a random vector , as (up to a multiplicative constant)
where is the marginal c.d.f. of the -th component of random vector .
Back to the empirical version, contrary to L2 distance (or more generally any Minkowski -distance with ) for which the distance is null if and only if , many cases may arise for the Gini distance function. Let be the rank vector of point , then
Consequently the Gini prametric is null if , a standard property of distance functions. This condition does not imply since the position of the values of (and ) depend on the rank within the features (columns) . The second case implies . This means that the Gini prametric is sensitive to the rank dependency of the features. When all features have the same rank vectors, then the distance between each and every pairs of points is null.
In general, the Gini prametric does not coincide with the Euclidean L2 distance nor the Manhattan L1 one, apart from specific cases. If a model based on needs to be trained again with a new observation , some rerankings of the points may occur implying new distances between the points. However, in some particular situations in which rank vectors are just translated by the same amount, the distance remains constant. With two points, and ,
But if there was a third point, , then the ranks of points and are rescaled, with respectively and , so that
Accordingly, a strategy for computing conditional ranks is necessary for train-test splits (this will be described in Section 4). The main properties of the Gini prametric are the following.
Proposition 3.2.
(Gini prametric)
Let such that , let a matrix of ones, and let be the set of positive integers:
if (Nullity)
if (Rank-Nullity)
(Non-Negativity)
if
(Linear Invariance)
if and
(Rank Invariance)
Proof.
Straightforward from the properties of the Gini covariance (Yitzhaki & Schechtman, 2013). ∎
Linear invariance postulates that if the elements in are increasing while their ranks remain constant, then the Gini prametric is invariant to such a translation. On the other hand, if a new point is included in , or if is transformed in a such a way that the ranks increase by the same amount, the Gini prametric remains invariant. Finally, it is non-negative and symmetric, therefore following the terminology employed by Arkhangelskii & Pontryagin (1990), it can be referred to as a Gini symmetric, such that is a Gini pramatric space.
Notice that if denotes descending ranks
3.4 The Generalized Gini Prametric
In a series of papers Schechtman & Yitzhaki (1987; 2003); Yitzhaki & Schechtman (2013) define a generalized version of the Gini Mean Difference:
with the decumulative distribution function (i.e., ). Yitzhaki & Schechtman (2013) show that the Gini covariance operator is a robust version of the standard covariance operator. The robustness to outliers allows Gini-based measures to be enough flexible thanks to the hyper-parameter . This flexibility yields robust estimators of regression parameters as well as correlation coefficients Yitzhaki & Lambert (2013). Additionally, in the case of for multivariate normal data, least squares regressions become a particular case of Gini regressions. Generalized Gini principal components analysis, as introduced by Charpentier et al. (2021), enables the influence of outliers to be minimized and helps find the best sub-space to fit the data according to the value of (for and multivariate normal data, the usual principal component analysis is recovered). Following Schechtman & Yitzhaki (2003), the use of decumulative ranks raised to the power of allows the Gini prametric to be rewritten as follows
Definition 3.3.
The Generalized Gini prametric is defined for all such that and are two given rows of (denoted “”):
for , where denotes the descending rank of element in .
This generalized Gini prametric has the same property as except Rank Invariance due to the hyper-parameter (if .
Proposition 3.4.
(Generalized Gini prametric)
The generalized Gini prametric
satisfies Nullity (both), Non-Negativity, Symmetry and Linear Invariance.
Proof.
Straightforward from the properties of the Gini covariance (Yitzhaki & Schechtman, 2013). ∎
The generalized version of offers more flexibility about the rank correlation between the features to be taken into account. As in the Gini-covariance, is referred to be the median value, which places as much weight on the lower part of the feature distributions as on the upper part. For , more weights are put to higher ranks, whereas for , more weights are put on lower ranks.
In order to interpret the correlations that lie behind the generalized Gini prametric, let us take centered variables and centered rank vectors , then:
The two first terms of the previous relation correspond to the two Gini covariances (up to ): with the rank of and conversely with the ranks of . The two last terms are (up to ) generalized Gini Mean Difference, which correspond to the intrinsic variability of and respectively. Everything happens as if the generalized Gini prametric measured on points belonging to different groups captures the two Gini covariances between groups in excess of the covariances within groups. In this respect, it could be included in the family of distances (F3) based on scalar products.
4 Gini KNN and Gini K-means
Algorithms implemented in a Gini prametric space may be used instead of those relying on standard L metric space for a given integer . We will discuss two algorithms in this section, Gini KNN (4.1) and Gini K-means (4.2).
4.1 Gini KNN
K-Nearest Neighbors (KNN) is a non-parametric algorithm used for classification or regression. After identifying the closest data points (neighbors) to a given input based on a distance metric (e.g. Gini prametric, Euclidean, Manhattan, etc.), the KNN consists in predicting the output based on the majority class (classification) or average value (regression) of these neighbors.
The Gini KNN relies on two hyper-parameters to learn during the training phase of the KNN: the number of neighbors and the value of the hyper-parameter . Since the approach is supervised there is no need to define ex ante the value, which can be deduced to be the best parameter that maximizes the F-measure or other pre-defined metrics (precision, recall, etc.). In the same time, the optimal value of may be deduced from the maximization of the same metrics. The prediction of the test points (regression task or classification task) depends on the computation of their conditional ranks, necessary for the computation of the generalized Gini prametric. The conditional ranks are ranks the test points would have if they were part of the training dataset. These are the last elements extracted from vector , with and the size of the testing data and the size of the whole dataset respectively.
Finally, the algorithm is convergent and its error rate is bounded.
Proposition 4.1.
(KNN convergence and error rate bound)
The KNN algorithm converges using the generalized Gini prametric , for all .
Let be the number of classes, the Bayes error rate and the asymptotic KNN error rate, then
Proof.
See Appendix A.1. ∎
4.2 Gini K-means
K-means is an unsupervised clustering algorithm that groups data points into clusters by minimizing the Euclidean distance (variance), or other distance metrics, within each cluster. It iteratively assigns each point to the nearest cluster centroid and updates the centroids by calculating the mean of the points within each cluster. The process is repeated until the centroids converge or a maximum number of iterations is reached.
The Gini K-means algorithm based on the generalized Gini prametric may be implemented if two requirements are met. First, convergence must be ensured for any given dataset and hyper-parameters .
Proposition 4.2.
The Gini K-means algorithm based on the generalized Gini prametric is convergent, for all .
Proof.
See Appendix A.2. ∎
Second, since convergence is always guaranteed, an optimal value may be derived. Setting training and testing data as , a grid search may be conducted as follows. The Gini K-means algorithm is trained on and the silhouette criterion (Rousseeuw, 1987) is computed for each value of (not exceeding 6, otherwise too much weight is placed on minimal values of ). The optimal hyper-parameter is the value that maximizes the silhouette score. Once is determined, the labels of the testing data can be inferred.
Once the hyper-parameter is obtained, Euclidean and Gini K-means may be compared, as depicted in Figure 1.
Finally, other clustering techniques relying on the Euclidean distance, close to K-means, may be adapted to the generalized Gini prametric, such as the agglomerative clustering (not detailed here). An illustration is depicted in Figure 2 on the iris dataset.111 See Github.
5 Experiments
K-means and KNN classifications have been performed adapting the scikit-learn library to the generalized Gini prametric. An additional illustration is also conducted with agglomerative clustering. These algorithms are compared on the basis of 12 distances issued from the 8 families (described in Table 1): L Minkowski distance (Euclidean, Minkowski, Manhattan), L1 distance (Canberra), Inner product distance (Cosine), Squared Chord distance (Hellinger distance), Squared L2 distance (Pearson ), Shannon entropy distance (Jensen-Shannon), Vicissitude distance (Vicis Symmetric), and other distances (Hassanat). The comparison is performed over 16 datasets of UCI repository222UCI datasets repository link., without normalization in order to observe the impact of outliers (if present) on the performance. Following the study of Alfeilat et al. (2019), the 16 datasets have been chosen with regard to the excellent performance of some distances robust to outliers, in particular the Hassanat distance. The datasets are described in Table 2. The detailed metrics of the Gini K-means and Gini KNN experiments, analyzed in the following subsections, are relegated in the Appendix.
| Name | #E | #F | #C | Data type | Min | Max |
| (A) Australian | 690 | 14 | 2 | Real | 0 | 100001 |
| (B) Banknote | 1372 | 4 | 2 | Real | -13.77 | 17.93 |
| (Bc) Breast-Cancer | 569 | 30 | 2 | Real, numeric | 1 | 13454352 |
| (I) Ionosphere | 351 | 34 | 2 | Real | -1 | 1 |
| (Ir) Iris | 150 | 4 | 3 | Real, numeric | 0.1 | 7.9 |
| (G) German | 1000 | 20 | 2 | Integer, Categorical | 0 | 184 |
| (Gl) Glass | 214 | 9 | 6 | Real | 0 | 75.41 |
| (He) Heart | 303 | 13 | 2 | Integer | 0 | 564 |
| (QS) QSAR | 1060 | 41 | 2 | Real | -5.256 | 147 |
| (S) Sonar | 208 | 60 | 2 | Real, Categorical | 0 | 1 |
| (V) Vehicle | 946 | 18 | 4 | Integer | 0 | 1018 |
| (Wi) Wine | 178 | 13 | 3 | Real, numeric | 813 | 1680 |
| (Ba) Balance | 625 | 4 | 3 | Positive Integer | 1 | 5 |
| (Ha) Haberman | 306 | 3 | 2 | Positive Integer | 0 | 83 |
| (W) Wholesale | 440 | 8 | 2 | Positive Integer | 1 | 112151 |
| (In) Indian Liver | 583 | 10 | 2 | Real, Integer | 0 | 297 |
5.1 KNN experiments
The generalized Gini prametric is implemented with the supervised KNN algorithm using the scikit-learn library. To compare the 12 distances listed above, precision and recall metrics are used on the same UCI datasets described in Table 2. To identify the optimal number of neighbors, is looped from to for each distance. A grid search is then performed to select the optimal hyper-parameter of the generalized Gini prametric that maximizes the precision. Then, recall and precision are computed and a 3-fold cross-validation is performed to validate the results (see Tables 16 and 12 in Appendix A, and in Appendix and B).
Without Noise. All datasets are treated as a task of classification (for regressions the output is discretized in order to compute the confusion matrix). Table 3 provides the precision rankings of the 12 KNN models evaluated across the 16 UCI datasets. The Gini KNN achieves the highest rank on 8 datasets, with an average rank of 3.12 (last column). The second-best KNN model, based on Hassanat distance, has an average rank of 4.06, with top positions on 5 datasets. These KNN models are followed by those relying on Hellinger distance (F4 family) and Gini (), with 4.62 and 5.12 average ranks, respectively.
| Distances | Ir | Wi | Bc | S | QS | V | A | Gl | I | B | G | He | Ba | Ha | W | In | Rank |
| Gini prametric | 1 | 6 | 4 | 1 | 1 | 1 | 4 | 1 | 4 | 1 | 1 | 4 | 3 | 10 | 1 | 7 | 3.12 |
| Gini prametric | 4 | 8 | 4 | 2 | 2 | 1 | 5 | 5 | 9 | 1 | 4 | 8 | 7 | 12 | 3 | 7 | 5.12 |
| Euclidean | 4 | 11 | 9 | 10 | 5 | 6 | 11 | 3 | 6 | 1 | 11 | 9 | 10 | 2 | 6 | 2 | 6.62 |
| Manhattan | 10 | 10 | 8 | 6 | 5 | 3 | 9 | 2 | 3 | 1 | 6 | 9 | 6 | 3 | 6 | 2 | 5.56 |
| Minkowski | 4 | 12 | 9 | 10 | 9 | 9 | 11 | 8 | 8 | 1 | 11 | 9 | 7 | 1 | 10 | 6 | 7.81 |
| Cosine | 1 | 9 | 9 | 6 | 9 | 4 | 10 | 6 | 6 | 7 | 10 | 12 | 2 | 3 | 12 | 12 | 7.38 |
| Hassanat | 11 | 3 | 1 | 2 | 9 | 4 | 2 | 3 | 1 | 1 | 1 | 1 | 10 | 3 | 6 | 7 | 4.06 |
| Canberra | 11 | 2 | 1 | 9 | 3 | 9 | 1 | 10 | 1 | 10 | 6 | 2 | 10 | 8 | 3 | 7 | 5.81 |
| Hellinger | 4 | 6 | 4 | 2 | 5 | 7 | 5 | 6 | 4 | 8 | 6 | 3 | 3 | 8 | 1 | 2 | 4.62 |
| Jensen-Shannon | 1 | 4 | 9 | 2 | 5 | 7 | 5 | 10 | 11 | 12 | 4 | 4 | 1 | 3 | 11 | 1 | 5.62 |
| Pearson Chi2 | 4 | 5 | 4 | 6 | 12 | 12 | 8 | 8 | 12 | 8 | 1 | 4 | 3 | 10 | 3 | 7 | 6.69 |
| Vicis Symmetric | 4 | 1 | 1 | 12 | 3 | 11 | 3 | 12 | 10 | 11 | 6 | 7 | 7 | 3 | 6 | 2 | 6.19 |
| Distances | Ir | Wi | Bc | S | QS | V | A | Gl | I | B | G | He | Ba | Ha | W | In | Rank |
| Gini prametric | 1 | 7 | 3 | 1 | 1 | 1 | 3 | 1 | 5 | 1 | 1 | 7 | 3 | 6 | 1 | 1 | 2.69 |
| Gini prametric | 2 | 8 | 3 | 2 | 1 | 1 | 7 | 3 | 9 | 1 | 4 | 8 | 3 | 12 | 3 | 1 | 4.25 |
| Euclidean | 2 | 11 | 9 | 9 | 3 | 6 | 11 | 2 | 5 | 1 | 11 | 10 | 6 | 6 | 7 | 1 | 6.25 |
| Manhattan | 8 | 10 | 8 | 4 | 3 | 3 | 9 | 3 | 3 | 1 | 10 | 9 | 10 | 1 | 5 | 1 | 5.5 |
| Minkowski | 2 | 12 | 9 | 9 | 9 | 9 | 12 | 11 | 8 | 1 | 11 | 12 | 6 | 1 | 10 | 6 | 8 |
| Cosine | 2 | 9 | 12 | 8 | 11 | 5 | 10 | 5 | 5 | 7 | 8 | 11 | 2 | 1 | 11 | 6 | 7.06 |
| Hassanat | 10 | 3 | 1 | 2 | 9 | 4 | 2 | 7 | 1 | 1 | 2 | 3 | 11 | 1 | 7 | 1 | 4.06 |
| Canberra | 10 | 2 | 3 | 9 | 3 | 9 | 1 | 7 | 2 | 10 | 4 | 1 | 12 | 6 | 7 | 6 | 5.75 |
| Hellinger | 2 | 5 | 3 | 4 | 3 | 7 | 6 | 5 | 3 | 8 | 8 | 2 | 6 | 1 | 1 | 6 | 4.38 |
| Jensen-Shannon | 8 | 4 | 9 | 4 | 3 | 7 | 5 | 10 | 12 | 12 | 4 | 3 | 1 | 6 | 11 | 12 | 6.94 |
| Pearson Chi2 | 2 | 5 | 3 | 4 | 12 | 12 | 7 | 9 | 11 | 8 | 2 | 3 | 3 | 6 | 3 | 6 | 6 |
| Vicis Symmetric | 10 | 1 | 1 | 12 | 3 | 11 | 3 | 12 | 10 | 11 | 4 | 3 | 6 | 11 | 5 | 11 | 7.12 |
The same results are obtained in terms of recall (Table 4). The Gini KNN achieves the top rank 9 times, while Hassanat achieves it 5 times. The standard Gini KNN () ranks third.
With 5% and 10% Noise. The experiment involves 5% and 10% noise into the data. Specifically, Gaussian noise with a mean of 0 and a standard deviation of 1 is added. This setup is designed to evaluate the robustness of the KNN algorithms under less favorable conditions, where the generalized Gini prametric is expected to outperform the other methods. In the first case, the Gini KNN achieves the highest average precision (11 top-ranked positions out of 16) and recall (10 top-ranked positions). In the 10% contamination case, the performance are much more variable. The generalized Gini prametric achieves only 6 best positions for precision and 5 for recall. Then, it comes Hassanat, Minkowski and Cosine (2 top positions for precision), and Hassanat again for recall (3 top positions), see Tables 2 to 6 in Appendix B.
5.2 K-means experiments
To ensure accurate performance measurement, a 5-fold cross-validation is performed with a fixed seed for each experiment. The number of clusters is set to i.e. the number of classes in Table 2. K-means are first initialized with the k-means++ algorithm Arthur & Vassilvitskii (2007), based on a sampling of the empirical probability distributions of the points (of the entire dataset) that accelerates convergence. All K-means models (issued from the 12 distances) are then initialized with the same optimized centroids. For Gini K-means, a grid search is performed over the hyper-parameter , selecting the value that maximized the precision score. Indeed, since the UCI datasets are supervised, computing the silhouette criterion is not required. After predicting the labels on the test data, the Kuhn-Munkres algorithm Kuhn (1955) is applied to align the predicted clusters with the true labels, as K-means does not assign labels corresponding to the actual classes. Finally, the average precision and recall are computed across all folds.
Without Noise. The generalized Gini prametric achieves the best performance thanks to its hyper-parameter. It is top-ranked on 5 datasets out of 16 (in terms of precision Table 26), with a mean rank of 3.5 (the best mean rank of 4.19 for recall Table 27). The second best method is the K-means based on the Hassanat distance in terms of precision and recall (mean rank of 5.06 and 4.75 respectively). The details about precision and recall are relegated in Tables 7-11 of Appendix B.
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini prametric | 4 | 1 | 1 | 1 | 1 | 5 | 6 | 3 | 4 | 6 | 3 | 1 | 8 | 6 | 2 | 4 | 3.50 |
| Gini prametric | 7 | 8 | 4 | 5 | 4 | 7 | 12 | 6 | 5 | 8 | 4 | 2 | 12 | 7 | 3 | 8 | 6.38 |
| Euclidean | 10 | 4 | 5 | 8 | 5 | 11 | 3 | 10 | 8 | 11 | 8 | 7 | 10 | 10 | 10 | 10 | 8.12 |
| Manhattan | 8 | 3 | 5 | 10 | 3 | 9 | 1 | 10 | 6 | 9 | 9 | 5 | 9 | 11 | 10 | 6 | 7.12 |
| Minkowski | 9 | 2 | 5 | 7 | 6 | 10 | 5 | 10 | 7 | 10 | 7 | 6 | 11 | 9 | 10 | 11 | 7.81 |
| Cosine | 6 | 6 | 10 | 3 | 2 | 8 | 2 | 8 | 12 | 7 | 5 | 8 | 6 | 12 | 5 | 7 | 6.69 |
| Canberra | 2 | 7 | 3 | 2 | 9 | 1 | 11 | 5 | 11 | 2 | 2 | 10 | 4 | 8 | 9 | 5 | 5.69 |
| Hellinger | 1 | 5 | 8 | 9 | 8 | 6 | 9 | 1 | 2 | 4 | 11 | 12 | 1 | 2 | 7 | 2 | 5.5 |
| Jensen-Shannon | 11 | 12 | 9 | 12 | 10 | 12 | 10 | 1 | 1 | 12 | 12 | 9 | 2 | 3 | 4 | 12 | 8.25 |
| Pearson Chi2 | 11 | 11 | 12 | 6 | 11 | 3 | 4 | 7 | 3 | 1 | 6 | 3 | 5 | 4 | 6 | 1 | 5.88 |
| Vicis Symmetric | 3 | 10 | 11 | 11 | 12 | 2 | 8 | 9 | 9 | 5 | 10 | 11 | 3 | 1 | 8 | 3 | 7.25 |
| Hassanat | 5 | 9 | 2 | 4 | 7 | 4 | 7 | 4 | 10 | 2 | 1 | 4 | 7 | 5 | 1 | 9 | 5.06 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini prametric | 4 | 1 | 2 | 2 | 2 | 9 | 7 | 7 | 4 | 6 | 2 | 1 | 6 | 6 | 4 | 4 | 4.19 |
| Gini prametric | 9 | 8 | 3 | 3 | 3 | 7 | 7 | 6 | 4 | 9 | 4 | 3 | 8 | 7 | 5 | 6 | 5.75 |
| Euclidean | 12 | 4 | 5 | 5 | 4 | 10 | 10 | 10 | 7 | 11 | 8 | 8 | 11 | 9 | 9 | 10 | 8.31 |
| Manhattan | 10 | 2 | 5 | 8 | 1 | 8 | 12 | 10 | 4 | 10 | 9 | 5 | 9 | 10 | 9 | 8 | 7.5 |
| Minkowski | 11 | 2 | 5 | 7 | 6 | 11 | 6 | 10 | 7 | 11 | 7 | 7 | 12 | 8 | 9 | 10 | 8.06 |
| Cosine | 6 | 5 | 11 | 4 | 5 | 6 | 2 | 4 | 12 | 8 | 5 | 9 | 4 | 12 | 2 | 7 | 6.38 |
| Canberra | 3 | 6 | 4 | 9 | 10 | 3 | 11 | 2 | 10 | 3 | 1 | 11 | 2 | 10 | 12 | 5 | 6.38 |
| Hellinger | 2 | 7 | 12 | 10 | 9 | 4 | 4 | 9 | 2 | 4 | 11 | 12 | 3 | 4 | 7 | 1 | 6.31 |
| Jensen-Shannon | 7 | 9 | 8 | 12 | 12 | 12 | 5 | 8 | 1 | 7 | 12 | 4 | 1 | 1 | 3 | 12 | 7.12 |
| Pearson Chi2 | 7 | 12 | 10 | 6 | 11 | 2 | 1 | 3 | 3 | 1 | 6 | 2 | 5 | 3 | 6 | 2 | 5 |
| Vicis Symmetric | 1 | 11 | 9 | 11 | 8 | 5 | 9 | 1 | 11 | 5 | 10 | 10 | 7 | 2 | 8 | 3 | 6.94 |
| Hassanat | 5 | 10 | 1 | 1 | 7 | 1 | 3 | 5 | 9 | 2 | 2 | 6 | 10 | 5 | 1 | 8 | 4.75 |
With 5% and 10% Noise. The same experiments are conducted for different levels of noise (5% and 10%) as in the KNN experiments, under the same conditions. The Gini K-means reaches the top position 3 times (out of 16 datasets) with a mean ranking of 3.75 (5% noise). The Hellinger distance is ranked 2nd, with more top-ranked positions (5 out of 16 datasets), but with a larger variability so that the mean rank is only 5 (like Hassanat). In the case of 10% noise, the generalized Gini prametric has only 2 top-ranked positions and a mean rank of 4.19. Again, Hellinger has much more top-ranked positions (5 out of 16) but a mean rank of 5.06, exhibiting a more important variability in the ranking (sometimes ranked 10 or 11). The results are presented in Appendix B (Tables 12-20).
Strategy to select for real applications. The performances described above have been computed with known labels for the test points. In real applications, once is fixed, a strategy is necessary to select the hyper-parameter. As described in Algorithm 1, a 5-fold cross-validation with a fixed seed is used to select the hyper-parameter that maximizes the silhouette score (mean across all folds). Tables 38 and 39 indicate that the generalized Gini prametric is ranked 4th in precision and 3rd in recall (for ). This experiment shows that maximizing the silhouette score is not the best method, but it allows the generalized Gini prametric to be ranked among the top 4 distances (Tables 21-24 in Appendix B).
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini prametric | 8 | 1 | 4 | 2 | 6 | 6 | 9 | 3 | 5 | 8 | 6 | 6 | 11 | 8 | 2 | 5 | 5.62 |
| Gini prametric | 6 | 8 | 3 | 5 | 3 | 7 | 12 | 6 | 4 | 7 | 3 | 1 | 12 | 6 | 3 | 8 | 5.88 |
| Euclidean | 10 | 4 | 5 | 8 | 4 | 11 | 3 | 10 | 8 | 11 | 8 | 7 | 9 | 10 | 10 | 10 | 8 |
| Manhattan | 7 | 3 | 5 | 10 | 2 | 9 | 1 | 10 | 6 | 9 | 9 | 4 | 8 | 11 | 10 | 6 | 6.88 |
| Minkowski | 9 | 2 | 5 | 7 | 5 | 10 | 5 | 10 | 7 | 10 | 7 | 5 | 10 | 9 | 10 | 11 | 7.62 |
| Cosine | 5 | 6 | 10 | 3 | 1 | 8 | 2 | 8 | 12 | 6 | 4 | 8 | 6 | 12 | 5 | 7 | 6.44 |
| Canberra | 2 | 7 | 2 | 1 | 9 | 1 | 11 | 5 | 11 | 2 | 2 | 10 | 4 | 7 | 9 | 4 | 5.44 |
| Hellinger | 1 | 5 | 8 | 9 | 8 | 5 | 8 | 1 | 2 | 4 | 11 | 12 | 1 | 2 | 7 | 2 | 5.38 |
| Jensen-Shannon | 11 | 12 | 9 | 12 | 10 | 12 | 10 | 1 | 1 | 12 | 12 | 9 | 2 | 3 | 4 | 12 | 8.25 |
| PearsonChi2 | 11 | 11 | 12 | 6 | 11 | 3 | 4 | 7 | 3 | 1 | 5 | 2 | 5 | 4 | 6 | 1 | 5.75 |
| VicisSymmetric1 | 3 | 10 | 11 | 11 | 12 | 2 | 7 | 9 | 9 | 5 | 10 | 11 | 3 | 1 | 8 | 3 | 7.19 |
| Hassanat | 4 | 9 | 1 | 4 | 7 | 4 | 6 | 4 | 10 | 2 | 1 | 3 | 7 | 5 | 1 | 9 | 4.81 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini prametric | 10 | 1 | 2 | 3 | 3 | 8 | 12 | 8 | 4 | 9 | 6 | 6 | 7 | 11 | 6 | 5 | 6.31 |
| Gini prametric | 8 | 8 | 3 | 2 | 2 | 7 | 7 | 6 | 4 | 8 | 3 | 2 | 8 | 6 | 4 | 6 | 5.25 |
| Euclidean | 12 | 4 | 5 | 5 | 4 | 10 | 9 | 10 | 7 | 11 | 8 | 8 | 11 | 8 | 9 | 10 | 8.19 |
| Manhattan | 9 | 2 | 5 | 8 | 1 | 8 | 11 | 10 | 4 | 10 | 9 | 4 | 9 | 9 | 9 | 8 | 7.25 |
| Minkowski | 11 | 2 | 5 | 7 | 6 | 11 | 6 | 10 | 7 | 11 | 7 | 7 | 12 | 7 | 9 | 10 | 8 |
| Cosine | 5 | 5 | 11 | 4 | 5 | 6 | 2 | 4 | 12 | 7 | 4 | 9 | 4 | 11 | 2 | 7 | 6.12 |
| Canberra | 3 | 6 | 4 | 9 | 10 | 3 | 10 | 2 | 10 | 3 | 1 | 11 | 2 | 9 | 12 | 4 | 6.19 |
| Hellinger | 2 | 7 | 12 | 10 | 9 | 4 | 4 | 9 | 2 | 4 | 11 | 12 | 3 | 4 | 7 | 1 | 6.31 |
| Jensen-Shannon | 6 | 9 | 8 | 12 | 12 | 12 | 5 | 7 | 1 | 6 | 12 | 3 | 1 | 1 | 3 | 12 | 6.88 |
| PearsonChi2 | 6 | 12 | 10 | 6 | 11 | 2 | 1 | 3 | 3 | 1 | 5 | 1 | 5 | 3 | 5 | 2 | 4.75 |
| VicisSymmetric1 | 1 | 11 | 9 | 11 | 8 | 5 | 8 | 1 | 11 | 5 | 10 | 10 | 6 | 2 | 8 | 3 | 6.81 |
| Hassanat | 4 | 10 | 1 | 1 | 7 | 1 | 3 | 5 | 9 | 2 | 2 | 5 | 10 | 5 | 1 | 8 | 4.62 |
5.3 Illustration on Agglomerative Hierarchical Clustering
For this additional experiment, noise has been directly added to the data using a standard normal distribution with a noise level of 10% to evaluate which distance metric demonstrates greater robustness under noisy conditions. A 3-fold cross-validation has been made and the linkage criterion has been fixed to ‘average’, i.e., the average distance between all points between clusters is used for all methods. The generalized Gini prametric is top-ranked with 7 top positions out of 16 datasets, and a mean rank of 1.94 (for precision). However, for the recall Ward’s technique is the best (7 top positions and an average rank of 2.62), see Appendix A.5.
6 Conclusion
The generalized Gini prametric, while not being a true distance, satisfies two invariance properties, making it robust to outliers. In particular, it is linearly invariant (similar to the Euclidean distance) and rank invariant. Accordingly, monotonic increasing transformations preserve ranks, ensuring that the Gini prametric exhibits small variations. Consequently, when the data contain outliers or measurement errors, the generalized Gini prametric helps achieve robust results, as shown on the 16 UCI datasets.
The experiments on the UCI datasets highlight the limitations of the generalized Gini prametric in unsupervised tasks. In particular, a strategy for selecting the hyper-parameter, such as using the silhouette score, must be chosen without prior knowledge of the quality of the results. Further grid searches for the hyper-parameter remain an open problem.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning with the Gini methodology. This method can help to refine standard clustering approach, particularly in the presence of outliers.
Appendix A Appendix
A.1 Proposition 4.1
The proof is exactly the same as the one of Cover (1967). The convergence is true for any type of metric used and the bound on the error rate does not change.
A.2 Proposition 4.2
Let be the centroid of group at iteration . If point goes from cluster at iteration to at iteration then:
The last inequality comes from the algorithm itself. Indeed, if , this is because . Now, we must prove that taking the sum over on the last expression implies:
Let us prove, for all points in a given cluster , that the mean is the unique point in implying a minimal distance, i.e.,
with . Taking the derivative of with respect to , and noting that for small variations of ranks remain constant, yields
Using the scalar product this is equivalent to:
Thus,
And finally,
Consequently, the unique centroid that minimizes the distance with all points in cluster is the arithmetic mean of points . Therefore, Eq.(1) is always true if .
A.3 Gini K-means without noise
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 3 3.52 | 0.0998 | 0.1 | 3 4.10 | 0.1913 | 0.1738 | 2 2.70 | 0.9234 | 0.9157 | 2 1.22 | 0.6339 | 0.5956 | 3 3.12 | 0.3741 | 0.2978 | 3 4.52 | 0.2946 | 0.2801 | 2 1.66 | 0.7693 | 0.6145 | 6 2.54 | 0.1492 | 0.0699 |
| Gini | 3 2 | 0.0828 | 0.1 | 3 2 | 0.1172 | 0.1005 | 2 2 | 0.9129 | 0.9052 | 2 2 | 0.5985 | 0.5618 | 3 2 | 0.3419 | 0.2644 | 3 2 | 0.2596 | 0.2517 | 2 2 | 0.6891 | 0.6101 | 6 2 | 0.0584 | 0.0327 |
| Euclidean | 3 - | 0.0757 | 0.0933 | 3 - | 0.0610 | 0.0502 | 2 - | 0.8802 | 0.8524 | 2 - | 0.5636 | 0.5373 | 3 - | 0.3520 | 0.2349 | 3 - | 0.2088 | 0.2269 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0388 | 0.0186 |
| Manhattan | 3 - | 0.0788 | 0.1 | 3 - | 0.0680 | 0.0613 | 2 - | 0.8782 | 0.8489 | 2 - | 0.5818 | 0.5518 | 3 - | 0.3531 | 0.2567 | 3 - | 0.2057 | 0.2234 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0628 | 0.0279 |
| Minkowski | 3 3 | 0.0774 | 0.0933 | 3 3 | 0.0643 | 0.0502 | 2 3 | 0.8812 | 0.8542 | 2 3 | 0.5784 | 0.5419 | 3 3 | 0.3498 | 0.2311 | 3 3 | 0.2100 | 0.2281 | 2 3 | 0.4873 | 0.5590 | 6 3 | 0.0359 | 0.0186 |
| Cosine | 3 - | 0.0188 | 0.02 | 3 - | 0.1370 | 0.1119 | 2 - | 0.9128 | 0.9051 | 2 - | 0.5627 | 0.5325 | 3 - | 0.3945 | 0.3055 | 3 - | 0.1984 | 0.2186 | 2 - | 0.6770 | 0.6725 | 6 - | 0.0599 | 0.0280 |
| Canberra | 3 - | 0.0386 | 0.0467 | 3 - | 0.3990 | 0.3430 | 2 - | 0.9237 | 0.9227 | 2 - | 0.5407 | 0.5225 | 3 - | 0.4449 | 0.3979 | 3 - | 0.2292 | 0.2234 | 2 - | 0.5208 | 0.5203 | 6 - | 0.1123 | 0.0421 |
| Hellinger | 3 - | 0.2906 | 0.2933 | 3 - | 0.2646 | 0.2740 | 2 - | 0.7313 | 0.7156 | 2 - | 0.3424 | 0.4901 | 3 - | 0.5476 | 0.3854 | 3 - | 0.3701 | 0.3275 | 2 - | 0.6701 | 0.5725 | 6 - | 0.1579 | 0.1728 |
| JensenShannon | 3 - | 0.4642 | 0.46 | 3 - | 0.0543 | 0.1352 | 2 - | 0.1272 | 0.3268 | 2 - | 0.5549 | 0.5531 | 3 - | 0.5241 | 0.5378 | 3 - | 0.3695 | 0.4196 | 2 - | 0.6830 | 0.6319 | 6 - | 0.0 | 0.0 |
| PearsonChi2 | 3 - | 0.2688 | 0.2867 | 3 - | 0.4505 | 0.4865 | 2 - | 0.8890 | 0.8683 | 2 - | 0.5973 | 0.5823 | 3 - | 0.3998 | 0.2990 | 3 - | 0.3644 | 0.3509 | 2 - | 0.6705 | 0.5913 | 6 - | 0.2025 | 0.1457 |
| VicisSymmetric1 | 3 - | 0.0667 | 0.0333 | 3 - | 0.2341 | 0.2037 | 2 - | 0.8656 | 0.8420 | 2 - | 0.4488 | 0.5296 | 3 - | 0.4766 | 0.2786 | 3 - | 0.4819 | 0.4031 | 2 - | 0.6502 | 0.5652 | 6 - | 0.1515 | 0.1260 |
| Hassanat | 3 - | 0.0497 | 0.06 | 3 - | 0.3990 | 0.3598 | 2 - | 0.9240 | 0.9157 | 2 - | 0.5834 | 0.5470 | 3 - | 0.3772 | 0.2490 | 3 - | 0.3563 | 0.2931 | 2 - | 0.7858 | 0.7739 | 6 - | 0.0553 | 0.0279 |
A.4 Gini KNN without noise
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 2 | 0.89 | 0.89 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 1 | 0.56 | 0.57 | 10 | 0.84 | 0.88 | 8 | 0.75 | 0.75 | 9 | 0.91 | 0.90 | 2 | 0.60 | 0.37 |
| Manhattan | 1 | 0.91 | 0.90 | 1 | 1.00 | 1.00 | 2 | 0.70 | 0.66 | 4 | 0.56 | 0.58 | 8 | 0.86 | 0.87 | 10 | 0.74 | 0.76 | 9 | 0.91 | 0.91 | 2 | 0.60 | 0.37 |
| Minkowski | 2 | 0.88 | 0.87 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 3 | 0.56 | 0.50 | 10 | 0.85 | 0.88 | 8 | 0.76 | 0.76 | 9 | 0.90 | 0.89 | 2 | 0.59 | 0.36 |
| Cosine | 2 | 0.89 | 0.89 | 1 | 0.99 | 0.99 | 8 | 0.69 | 0.69 | 6 | 0.52 | 0.55 | 3 | 0.93 | 0.90 | 2 | 0.74 | 0.76 | 10 | 0.88 | 0.88 | 1 | 0.55 | 0.36 |
| Hassanat | 2 | 0.92 | 0.92 | 1 | 1.00 | 1.00 | 10 | 0.73 | 0.71 | 3 | 0.67 | 0.67 | 3 | 0.84 | 0.82 | 4 | 0.74 | 0.76 | 10 | 0.91 | 0.90 | 2 | 0.58 | 0.37 |
| Canberra | 6 | 0.92 | 0.91 | 1 | 0.93 | 0.93 | 3 | 0.70 | 0.70 | 9 | 0.66 | 0.71 | 3 | 0.84 | 0.79 | 5 | 0.73 | 0.75 | 6 | 0.92 | 0.90 | 2 | 0.58 | 0.36 |
| Hellinger | 2 | 0.90 | 0.90 | 1 | 0.95 | 0.95 | 8 | 0.70 | 0.69 | 8 | 0.65 | 0.69 | 8 | 0.87 | 0.88 | 7 | 0.73 | 0.76 | 5 | 0.93 | 0.93 | 2 | 0.60 | 0.36 |
| Jensen-Shannon | 2 | 0.78 | 0.40 | 1 | 0.89 | 0.89 | 10 | 0.71 | 0.70 | 10 | 0.63 | 0.67 | 1 | 0.94 | 0.94 | 7 | 0.74 | 0.75 | 9 | 0.89 | 0.88 | 2 | 0.72 | 0.06 |
| PearsonChi2 | 1 | 0.41 | 0.64 | 1 | 0.95 | 0.95 | 8 | 0.73 | 0.71 | 7 | 0.63 | 0.67 | 8 | 0.87 | 0.89 | 8 | 0.72 | 0.75 | 7 | 0.92 | 0.92 | 2 | 0.58 | 0.36 |
| VicisSymmetric1 | 7 | 0.81 | 0.76 | 1 | 0.92 | 0.92 | 8 | 0.70 | 0.70 | 10 | 0.62 | 0.67 | 8 | 0.85 | 0.88 | 5 | 0.74 | 0.71 | 8 | 0.91 | 0.91 | 2 | 0.60 | 0.35 |
| Gini | 2 2 | 0.85 | 0.84 | 1 2 | 1.00 | 1.00 | 6 2 | 0.71 | 0.70 | 1 2 | 0.61 | 0.61 | 10 2 | 0.85 | 0.89 | 5 2 | 0.71 | 0.69 | 5 | 0.92 | 0.92 | 2 | 0.58 | 0.37 |
| Generalized Gini | 2 4.89 | 0.90 | 0.89 | 1 1.96 | 1.00 | 1.00 | 7 3.35 | 0.73 | 0.73 | 1 2.12 | 0.63 | 0.63 | 7 5.09 | 0.87 | 0.89 | 5 5.09 | 0.72 | 0.75 | 8 5.09 | 0.93 | 0.93 | 2 2.67 | 0.58 | 0.37 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 5 - | 0.97 | 0.97 | 1 - | 0.71 | 0.71 | 10 - | 0.93 | 0.93 | 1 - | 0.84 | 0.83 | 6 - | 0.64 | 0.65 | 1 - | 0.87 | 0.87 | 6 - | 0.70 | 0.69 | 3 - | 0.74 | 0.72 |
| Manhattan | 1 - | 0.96 | 0.96 | 1 - | 0.77 | 0.77 | 6 - | 0.94 | 0.94 | 1 - | 0.86 | 0.86 | 9 - | 0.64 | 0.65 | 1 - | 0.90 | 0.90 | 8 - | 0.75 | 0.73 | 1 - | 0.75 | 0.71 |
| Minkowski | 5 - | 0.97 | 0.97 | 1 - | 0.68 | 0.69 | 10 - | 0.93 | 0.93 | 3 - | 0.84 | 0.83 | 8 - | 0.63 | 0.64 | 1 - | 0.85 | 0.85 | 10 - | 0.70 | 0.67 | 2 - | 0.68 | 0.65 |
| Cosine | 7 - | 0.98 | 0.97 | 1 - | 0.84 | 0.84 | 7 - | 0.93 | 0.92 | 1 - | 0.86 | 0.85 | 9 - | 0.63 | 0.63 | 1 - | 0.89 | 0.88 | 9 - | 0.72 | 0.72 | 8 - | 0.69 | 0.69 |
| Hassanat | 10 - | 0.95 | 0.95 | 1 - | 0.96 | 0.95 | 7 - | 0.96 | 0.96 | 1 - | 0.87 | 0.87 | 8 - | 0.63 | 0.64 | 7 - | 0.89 | 0.89 | 8 - | 0.86 | 0.84 | 4 - | 0.74 | 0.68 |
| Canberra | 9 - | 0.95 | 0.95 | 1 - | 0.97 | 0.97 | 6 - | 0.96 | 0.95 | 1 - | 0.85 | 0.83 | 7 - | 0.65 | 0.65 | 1 - | 0.85 | 0.85 | 6 - | 0.87 | 0.86 | 2 - | 0.67 | 0.68 |
| Hellinger | 2 - | 0.97 | 0.97 | 4 - | 0.91 | 0.92 | 4 - | 0.95 | 0.95 | 1 - | 0.87 | 0.86 | 6 - | 0.64 | 0.65 | 1 - | 0.86 | 0.86 | 2 - | 0.77 | 0.76 | 1 - | 0.69 | 0.69 |
| Gini | 4 2 | 0.97 | 0.97 | 5 2 | 0.90 | 0.89 | 5 2 | 0.95 | 0.95 | 1 2 | 0.87 | 0.87 | 7 - | 0.66 | 0.67 | 1 2 | 0.93 | 0.93 | 6 2 | 0.77 | 0.75 | 3 2 | 0.73 | 0.71 |
| Generalized Gini | 7 1.1 | 0.98 | 0.98 | 3 4.45 | 0.91 | 0.91 | 4 3.11 | 0.95 | 0.95 | 1 2.6 | 0.88 | 0.88 | 7 2.39 | 0.67 | 0.67 | 1 3.48 | 0.93 | 0.93 | 5 3.45 | 0.79 | 0.79 | 3 1.37 | 0.77 | 0.74 |
| JensenShannon | 3 - | 0.98 | 0.96 | 10 - | 0.95 | 0.94 | 3 - | 0.93 | 0.93 | 1 - | 0.87 | 0.86 | 9 - | 0.64 | 0.65 | 9 - | 0.86 | 0.86 | 9 - | 0.77 | 0.78 | 1 - | 0.67 | 0.66 |
| PearsonChi2 | 2 - | 0.97 | 0.97 | 4 - | 0.92 | 0.92 | 4 - | 0.95 | 0.95 | 1 - | 0.86 | 0.86 | 2 - | 0.57 | 0.59 | 7 - | 0.76 | 0.75 | 7 - | 0.76 | 0.75 | 1 - | 0.68 | 0.67 |
| VicisSymmetric1 | 5 - | 0.97 | 0.95 | 1 - | 0.98 | 0.98 | 5 - | 0.96 | 0.96 | 1 - | 0.79 | 0.78 | 9 - | 0.65 | 0.65 | 5 - | 0.78 | 0.78 | 8 - | 0.81 | 0.79 | 3 - | 0.65 | 0.62 |
A.5 Agglomerative clustering with 10% noise
| Iris | Wine | Breast Cancer | Sonar | Australian | Glass | QSAR | Vehicle | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 3 2.37 | 0.8990 | 0.8867 | 3 4.27 | 0.7864 | 0.7187 | 2 2.50 | 0.9161 | 0.9050 | 2 2.54 | 0.4766 | 0.5336 | 3 2.22 | 0.7647 | 0.5957 | 6 1.20 | 0.5835 | 0.4485 | 3 1.36 | 0.6227 | 0.5982 | 4 4.95 | 0.6631 | 0.6028 |
| Gini | 3 - | 0.8808 | 0.8533 | 3 - | 0.5433 | 0.6575 | 2 - | 0.8697 | 0.8312 | 2 - | 0.3509 | 0.5239 | 3 - | 0.7153 | 0.5826 | 6 - | 0.5573 | 0.4765 | 3 - | 0.3992 | 0.5738 | 4 - | 0.4779 | 0.5035 |
| Euclidean | 3 - | 0.8798 | 0.8133 | 3 - | 0.5077 | 0.6345 | 2 - | 0.7767 | 0.6520 | 2 - | 0.4187 | 0.5239 | 3 - | 0.4434 | 0.5565 | 6 - | 0.4843 | 0.4068 | 3 - | 0.5287 | 0.5982 | 4 - | 0.6847 | 0.5330 |
| Euclidean Ward | 3 - | 0.8773 | 0.8600 | 3 - | 0.7941 | 0.7797 | 2 - | 0.8872 | 0.8632 | 2 - | 0.2947 | 0.5429 | 3 - | 0.7664 | 0.6174 | 6 - | 0.6302 | 0.4648 | 3 - | 0.5360 | 0.5946 | 4 - | 0.6592 | 0.5745 |
| Manhattan | 3 - | 0.8945 | 0.8333 | 3 - | 0.5104 | 0.6511 | 2 - | 0.7767 | 0.6520 | 2 - | 0.3741 | 0.5193 | 3 - | 0.4434 | 0.5565 | 6 - | 0.4024 | 0.4068 | 3 - | 0.5540 | 0.5969 | 4 - | 0.6878 | 0.5366 |
| Minkowski | 3 3 | 0.7308 | 0.6933 | 3 3 | 0.5104 | 0.6511 | 2 3 | 0.7767 | 0.6520 | 2 3 | 0.2848 | 0.5336 | 3 3 | 0.4434 | 0.5565 | 6 3 | 0.4319 | 0.4673 | 3 3 | 0.5657 | 0.6008 | 4 3 | 0.6028 | 0.5485 |
| Cosine | 3 - | 0.6619 | 0.7667 | 3 - | 0.4529 | 0.5901 | 2 - | 0.7831 | 0.6660 | 2 - | 0.2848 | 0.5336 | 3 - | 0.4771 | 0.5580 | 6 - | 0.4272 | 0.4626 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4863 | 0.5047 |
| Canberra | 3 - | 0.4870 | 0.6533 | 3 - | 0.4728 | 0.6518 | 2 - | 0.3940 | 0.6274 | 2 - | 0.4386 | 0.5384 | 3 - | 0.3095 | 0.5551 | 6 - | 0.3332 | 0.4439 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4879 | 0.4941 |
| Hellinger | 3 - | 0.8781 | 0.7667 | 3 - | 0.5793 | 0.6513 | 2 - | 0.7755 | 0.6484 | 2 - | 0.2848 | 0.5336 | 3 - | 0.7562 | 0.5609 | 6 - | 0.4082 | 0.4859 | 3 - | 0.4302 | 0.6033 | 4 - | 0.4774 | 0.5130 |
| JensenShannon | 3 - | 0.4870 | 0.6533 | 3 - | 0.2890 | 0.4878 | 2 - | 0.3940 | 0.6274 | 2 - | 0.2848 | 0.5336 | 3 - | 0.5377 | 0.6101 | 6 - | 0.1265 | 0.3552 | 3 - | 0.3491 | 0.5905 | 4 - | 0.2588 | 0.5071 |
| PearsonChi2 | 3 - | 0.8781 | 0.7667 | 3 - | 0.6765 | 0.6964 | 2 - | 0.6517 | 0.6379 | 2 - | 0.3484 | 0.5143 | 3 - | 0.7562 | 0.5609 | 6 - | 0.4082 | 0.4859 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4757 | 0.5213 |
| VicisSymmetric1 | 3 - | 0.4870 | 0.6533 | 3 - | 0.3309 | 0.4045 | 2 - | 0.3940 | 0.6274 | 2 - | 0.3588 | 0.5193 | 3 - | 0.5318 | 0.6014 | 6 - | 0.4604 | 0.4623 | 3 - | 0.3491 | 0.5905 | 4 - | 0.2984 | 0.5024 |
| Hassanat | 3 - | 0.7904 | 0.7867 | 3 - | 0.6191 | 0.7365 | 2 - | 0.8660 | 0.8314 | 2 - | 0.3458 | 0.5288 | 3 - | 0.6869 | 0.5768 | 6 - | 0.5249 | 0.4907 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4873 | 0.4917 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 7 | 1 | 2 | 2 | 2 | 1 | 1 | 2 | 1 | 2 | 1 | 1 | 2 | 2 | 1 | 3 | 1.94 |
| Gini | 7 | 12 | 3 | 5 | 10 | 4 | 13 | 2 | 3 | 6 | 3 | 6 | 5 | 3 | 7 | 9 | 6.12 |
| Euclidean | 7 | 3 | 10 | 9 | 11 | 4 | 7 | 2 | 4 | 9 | 6 | 3 | 10 | 5 | 5 | 2 | 6.06 |
| Euclidean Ward | 2 | 6 | 1 | 13 | 1 | 3 | 2 | 1 | 7 | 1 | 2 | 9 | 1 | 1 | 4 | 4 | 3.62 |
| Manhattan | 6 | 2 | 4 | 9 | 8 | 4 | 7 | 2 | 2 | 7 | 6 | 4 | 10 | 11 | 3 | 1 | 5.38 |
| Minkowski | 7 | 4 | 5 | 9 | 13 | 4 | 7 | 2 | 9 | 7 | 6 | 10 | 10 | 7 | 2 | 5 | 6.69 |
| Cosine | 7 | 7 | 6 | 8 | 9 | 4 | 3 | 2 | 10 | 11 | 5 | 10 | 9 | 8 | 8 | 8 | 7.19 |
| Canberra | 5 | 8 | 12 | 3 | 3 | 4 | 7 | 2 | 11 | 10 | 11 | 2 | 13 | 12 | 8 | 6 | 7.31 |
| Hellinger | 7 | 5 | 7 | 7 | 7 | 4 | 5 | 2 | 5 | 5 | 9 | 10 | 3 | 9 | 6 | 10 | 6.31 |
| JensenShannon | 7 | 13 | 12 | 6 | 4 | 4 | 4 | 2 | 11 | 13 | 11 | 10 | 7 | 13 | 8 | 13 | 8.62 |
| PearsonChi2 | 3 | 10 | 8 | 9 | 12 | 2 | 6 | 2 | 5 | 3 | 10 | 7 | 3 | 9 | 8 | 11 | 6.75 |
| VicisSymmetric1 | 4 | 9 | 11 | 4 | 6 | 4 | 7 | 2 | 11 | 12 | 11 | 5 | 8 | 6 | 8 | 12 | 7.5 |
| Hassanat | 1 | 11 | 9 | 1 | 5 | 4 | 7 | 2 | 8 | 4 | 4 | 8 | 6 | 4 | 8 | 7 | 5.56 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 7 | 1 | 2 | 3 | 8 | 3 | 1 | 2 | 1 | 3 | 1 | 3 | 4 | 9 | 3 | 1 | 3.25 |
| Gini | 7 | 12 | 4 | 13 | 1 | 3 | 13 | 2 | 3 | 5 | 4 | 9 | 5 | 4 | 13 | 10 | 6.75 |
| Euclidean | 7 | 4 | 6 | 8 | 11 | 3 | 7 | 2 | 5 | 10 | 6 | 9 | 10 | 11 | 3 | 5 | 6.69 |
| Euclidean Ward | 1 | 6 | 1 | 4 | 5 | 1 | 2 | 1 | 2 | 1 | 2 | 1 | 1 | 6 | 6 | 2 | 2.62 |
| Manhattan | 6 | 2 | 3 | 8 | 3 | 3 | 7 | 2 | 4 | 8 | 6 | 11 | 10 | 11 | 5 | 4 | 5.81 |
| Minkowski | 7 | 3 | 5 | 8 | 13 | 3 | 7 | 2 | 10 | 8 | 6 | 3 | 10 | 5 | 2 | 3 | 5.94 |
| Cosine | 7 | 7 | 9 | 5 | 6 | 3 | 3 | 2 | 7 | 11 | 5 | 3 | 9 | 7 | 7 | 9 | 6.25 |
| Canberra | 3 | 10 | 9 | 2 | 4 | 3 | 7 | 2 | 11 | 6 | 11 | 2 | 13 | 10 | 7 | 12 | 7 |
| Hellinger | 7 | 5 | 9 | 5 | 2 | 3 | 5 | 2 | 7 | 7 | 9 | 3 | 7 | 2 | 1 | 7 | 5.06 |
| JensenShannon | 7 | 11 | 9 | 5 | 9 | 3 | 4 | 2 | 11 | 12 | 11 | 3 | 2 | 13 | 7 | 8 | 7.31 |
| PearsonChi2 | 5 | 13 | 13 | 8 | 12 | 2 | 6 | 2 | 7 | 4 | 10 | 13 | 7 | 2 | 7 | 6 | 7.31 |
| VicisSymmetric1 | 2 | 9 | 8 | 8 | 7 | 3 | 7 | 2 | 11 | 13 | 11 | 11 | 3 | 8 | 7 | 11 | 7.56 |
| Hassanat | 3 | 8 | 7 | 1 | 10 | 3 | 7 | 2 | 6 | 2 | 3 | 8 | 6 | 1 | 7 | 13 | 5.44 |
Appendix B Experiments
In what follows, the results of the Minkowski distance are given for .
B.1 KNN experiments
B.1.1 Without noise
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 2 | 0.89 | 0.89 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 1 | 0.56 | 0.57 | 10 | 0.84 | 0.88 | 8 | 0.75 | 0.75 | 9 | 0.91 | 0.90 | 2 | 0.60 | 0.37 |
| Manhattan | 1 | 0.91 | 0.90 | 1 | 1.00 | 1.00 | 2 | 0.70 | 0.66 | 4 | 0.56 | 0.58 | 8 | 0.86 | 0.87 | 10 | 0.74 | 0.76 | 9 | 0.91 | 0.91 | 2 | 0.60 | 0.37 |
| Minkowski | 2 | 0.88 | 0.87 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 3 | 0.56 | 0.50 | 10 | 0.85 | 0.88 | 8 | 0.76 | 0.76 | 9 | 0.90 | 0.89 | 2 | 0.59 | 0.36 |
| Cosine | 2 | 0.89 | 0.89 | 1 | 0.99 | 0.99 | 8 | 0.69 | 0.69 | 6 | 0.52 | 0.55 | 3 | 0.93 | 0.90 | 2 | 0.74 | 0.76 | 10 | 0.88 | 0.88 | 1 | 0.55 | 0.36 |
| Hassanat | 2 | 0.92 | 0.92 | 1 | 1.00 | 1.00 | 10 | 0.73 | 0.71 | 3 | 0.67 | 0.67 | 3 | 0.84 | 0.82 | 4 | 0.74 | 0.76 | 10 | 0.91 | 0.90 | 2 | 0.58 | 0.37 |
| Canberra | 6 | 0.92 | 0.91 | 1 | 0.93 | 0.93 | 3 | 0.70 | 0.70 | 9 | 0.66 | 0.71 | 3 | 0.84 | 0.79 | 5 | 0.73 | 0.75 | 6 | 0.92 | 0.90 | 2 | 0.58 | 0.36 |
| Hellinger | 2 | 0.90 | 0.90 | 1 | 0.95 | 0.95 | 8 | 0.70 | 0.69 | 8 | 0.65 | 0.69 | 8 | 0.87 | 0.88 | 7 | 0.73 | 0.76 | 5 | 0.93 | 0.93 | 2 | 0.60 | 0.36 |
| Jensen-Shannon | 2 | 0.78 | 0.40 | 1 | 0.89 | 0.89 | 10 | 0.71 | 0.70 | 10 | 0.63 | 0.67 | 1 | 0.94 | 0.94 | 7 | 0.74 | 0.75 | 9 | 0.89 | 0.88 | 2 | 0.72 | 0.06 |
| PearsonChi2 | 1 | 0.41 | 0.64 | 1 | 0.95 | 0.95 | 8 | 0.73 | 0.71 | 7 | 0.63 | 0.67 | 8 | 0.87 | 0.89 | 8 | 0.72 | 0.75 | 7 | 0.92 | 0.92 | 2 | 0.58 | 0.36 |
| Vicis Symmetric | 7 | 0.81 | 0.76 | 1 | 0.92 | 0.92 | 8 | 0.70 | 0.70 | 10 | 0.62 | 0.67 | 8 | 0.85 | 0.88 | 5 | 0.74 | 0.71 | 8 | 0.91 | 0.91 | 2 | 0.60 | 0.35 |
| Gini | 2 2 | 0.85 | 0.84 | 1 2 | 1.00 | 1.00 | 6 2 | 0.71 | 0.70 | 1 2 | 0.61 | 0.61 | 10 2 | 0.85 | 0.89 | 5 2 | 0.71 | 0.69 | 5 | 0.92 | 0.92 | 2 | 0.58 | 0.37 |
| Generalized Gini | 2 4.89 | 0.90 | 0.89 | 1 1.96 | 1.00 | 1.00 | 7 3.35 | 0.73 | 0.73 | 1 2.12 | 0.63 | 0.63 | 7 5.09 | 0.87 | 0.89 | 5 5.09 | 0.72 | 0.75 | 8 5.09 | 0.93 | 0.93 | 2 2.67 | 0.58 | 0.37 |
B.1.2 With noise at level 5%
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 2 | 0.885 | 0.883 | 2 | 1.000 | 1.000 | 2 | 0.683 | 0.647 | 1 | 0.563 | 0.568 | 9 | 0.840 | 0.883 | 8 | 0.747 | 0.765 | 1 | 0.905 | 0.902 | 1 | 0.682 | 0.668 |
| Manhattan | 2 | 0.901 | 0.900 | 1 | 1.000 | 1.000 | 2 | 0.701 | 0.663 | 4 | 0.568 | 0.594 | 8 | 0.830 | 0.875 | 10 | 0.745 | 0.761 | 9 | 0.911 | 0.909 | 1 | 0.685 | 0.668 |
| Minkowski | 2 | 0.889 | 0.886 | 2 | 1.000 | 1.000 | 4 | 0.687 | 0.664 | 1 | 0.545 | 0.554 | 8 | 0.838 | 0.869 | 8 | 0.750 | 0.761 | 9 | 0.895 | 0.893 | 1 | 0.685 | 0.672 |
| Cosine | 2 | 0.907 | 0.903 | 1 | 0.998 | 0.998 | 8 | 0.696 | 0.690 | 6 | 0.541 | 0.564 | 3 | 0.916 | 0.907 | 9 | 0.738 | 0.754 | 10 | 0.881 | 0.877 | 1 | 0.670 | 0.668 |
| Hassanat | 2 | 0.904 | 0.903 | 4 | 0.999 | 0.999 | 10 | 0.722 | 0.713 | 9 | 0.657 | 0.703 | 8 | 0.821 | 0.872 | 4 | 0.738 | 0.764 | 8 | 0.907 | 0.902 | 9 | 0.698 | 0.711 |
| Canberra | 6 | 0.900 | 0.889 | 1 | 0.993 | 0.993 | 8 | 0.698 | 0.696 | 10 | 0.680 | 0.723 | 3 | 0.824 | 0.829 | 4 | 0.741 | 0.764 | 6 | 0.916 | 0.914 | 8 | 0.674 | 0.713 |
| Hellinger | 2 | 0.895 | 0.895 | 1 | 0.945 | 0.945 | 10 | 0.713 | 0.703 | 7 | 0.633 | 0.673 | 8 | 0.852 | 0.873 | 5 | 0.739 | 0.751 | 5 | 0.928 | 0.925 | 1 | 0.701 | 0.694 |
| Gini | 2 | 0.854 | 0.849 | 2 | 0.999 | 0.999 | 10 | 0.711 | 0.703 | 2 | 0.602 | 0.580 | 9 | 0.844 | 0.872 | 7 | 0.704 | 0.735 | 5 | 0.920 | 0.916 | 1 | 0.682 | 0.675 |
| Generalized Gini | 2 3.8 | 0.898 | 0.895 | 2 2.14 | 1.000 | 1.000 | 10 3.0 | 0.722 | 0.717 | 4 3.03 | 0.656 | 0.647 | 3 4.53 | 0.853 | 0.850 | 6 1.47 | 0.758 | 0.768 | 8 3.56 | 0.930 | 0.927 | 9 5.01 | 0.703 | 0.715 |
| JensenShannon | 2 | 0.772 | 0.373 | 4 | 0.747 | 0.537 | 2 | 0.603 | 0.557 | 1 | 0.766 | 0.261 | 1 | 0.743 | 0.835 | 7 | 0.745 | 0.765 | 9 | 0.885 | 0.875 | 2 | 0.687 | 0.718 |
| PearsonChi2 | 4 | 0.902 | 0.897 | 1 | 0.945 | 0.945 | 8 | 0.719 | 0.708 | 3 | 0.622 | 0.607 | 8 | 0.857 | 0.876 | 3 | 0.739 | 0.751 | 7 | 0.923 | 0.920 | 1 | 0.691 | 0.686 |
| VicisSymmetric1 | 2 | 0.875 | 0.872 | 1 | 0.985 | 0.985 | 4 | 0.651 | 0.633 | 3 | 0.654 | 0.647 | 8 | 0.859 | 0.850 | 5 | 0.733 | 0.752 | 6 | 0.923 | 0.920 | 10 | 0.682 | 0.710 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 9 | 0.982 | 0.980 | 1 | 0.706 | 0.707 | 10 | 0.935 | 0.935 | 1 | 0.834 | 0.832 | 6 | 0.644 | 0.653 | 1 | 0.865 | 0.863 | 9 | 0.711 | 0.703 | 1 | 0.702 | 0.678 |
| Manhattan | 10 | 0.976 | 0.973 | 1 | 0.776 | 0.775 | 10 | 0.941 | 0.940 | 3 | 0.845 | 0.845 | 9 | 0.633 | 0.643 | 1 | 0.900 | 0.898 | 8 | 0.745 | 0.727 | 1 | 0.741 | 0.701 |
| Minkowski | 5 | 0.980 | 0.980 | 1 | 0.682 | 0.685 | 10 | 0.933 | 0.933 | 3 | 0.805 | 0.803 | 6 3 | 0.646 | 0.653 | 1 | 0.849 | 0.846 | 10 | 0.703 | 0.696 | 1 | 0.707 | 0.696 |
| Cosine | 3 | 0.982 | 0.980 | 1 | 0.828 | 0.826 | 7 | 0.927 | 0.924 | 1 | 0.851 | 0.846 | 7 | 0.618 | 0.623 | 1 | 0.886 | 0.884 | 9 | 0.716 | 0.716 | 1 | 0.720 | 0.710 |
| Hassanat | 4 | 0.968 | 0.967 | 7 | 0.960 | 0.955 | 6 | 0.955 | 0.954 | 1 | 0.823 | 0.817 | 10 | 0.642 | 0.651 | 7 | 0.886 | 0.884 | 8 | 0.851 | 0.844 | 4 | 0.681 | 0.645 |
| Canberra | 6 | 0.968 | 0.967 | 7 | 0.958 | 0.955 | 9 | 0.939 | 0.938 | 10 | 0.793 | 0.658 | 9 | 0.641 | 0.625 | 10 | 0.869 | 0.869 | 10 | 0.863 | 0.862 | 1 | 0.514 | 0.509 |
| Hellinger | 4 | 0.981 | 0.980 | 4 | 0.918 | 0.915 | 4 | 0.946 | 0.945 | 1 | 0.853 | 0.846 | 9 | 0.657 | 0.664 | 1 | 0.863 | 0.861 | 5 | 0.767 | 0.763 | 3 | 0.664 | 0.668 |
| Gini | 5 | 0.974 | 0.973 | 5 | 0.898 | 0.893 | 3 | 0.944 | 0.943 | 1 | 0.857 | 0.856 | 7 | 0.660 | 0.668 | 3 | 0.925 | 0.924 | 4 | 0.766 | 0.746 | 4 | 0.736 | 0.696 |
| Generalized Gini | 9 1.1 | 0.987 | 0.987 | 1 4.52 | 0.918 | 0.916 | 4 3.07 | 0.950 | 0.949 | 1 2.93 | 0.872 | 0.870 | 7 1.89 | 0.672 | 0.678 | 1 3.51 | 0.937 | 0.936 | 4 4.30 | 0.796 | 0.777 | 3 2.40 | 0.752 | 0.705 |
| JensenShannon | 1 | 0.982 | 0.980 | 4 | 0.946 | 0.944 | 4 | 0.716 | 0.647 | 4 | 0.285 | 0.534 | 2 | 0.584 | 0.607 | 3 | 0.848 | 0.841 | 9 | 0.708 | 0.643 | 1 | 0.554 | 0.393 |
| PearsonChi2 | 4 | 0.981 | 0.980 | 4 | 0.918 | 0.916 | 4 | 0.946 | 0.945 | 1 | 0.847 | 0.841 | 9 | 0.667 | 0.671 | 4 | 0.865 | 0.864 | 7 | 0.760 | 0.755 | 4 | 0.668 | 0.654 |
| VicisSymmetric1 | 7 | 0.974 | 0.973 | 3 | 0.964 | 0.961 | 10 | 0.887 | 0.886 | 10 | 0.753 | 0.697 | 9 | 0.608 | 0.614 | 3 | 0.766 | 0.768 | 6 | 0.834 | 0.829 | 1 | 0.554 | 0.561 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Euclidean | 9 | 1 | 10 | 10 | 7 | 3 | 9 | 8 | 2 | 11 | 8 | 7 | 6 | 7 | 10 | 6 | 7.12 |
| Manhattan | 4 | 1 | 6 | 9 | 9 | 4 | 7 | 6 | 8 | 10 | 6 | 6 | 9 | 3 | 8 | 2 | 6.12 |
| Minkowski | 8 | 1 | 9 | 11 | 8 | 2 | 10 | 6 | 7 | 12 | 9 | 9 | 5 | 10 | 12 | 5 | 7.75 |
| Cosine | 1 | 7 | 8 | 12 | 1 | 9 | 12 | 12 | 2 | 9 | 10 | 4 | 10 | 4 | 9 | 4 | 7.12 |
| Hassanat | 2 | 5 | 1 | 3 | 11 | 9 | 8 | 3 | 11 | 2 | 1 | 8 | 7 | 4 | 2 | 7 | 5.25 |
| Canberra | 5 | 8 | 7 | 2 | 10 | 6 | 6 | 11 | 11 | 3 | 7 | 10 | 8 | 6 | 1 | 12 | 7.06 |
| Hellinger | 7 | 10 | 4 | 6 | 5 | 7 | 2 | 2 | 5 | 5 | 3 | 3 | 4 | 9 | 5 | 9 | 5.38 |
| Gini | 11 | 5 | 5 | 8 | 6 | 12 | 5 | 8 | 9 | 8 | 5 | 2 | 3 | 2 | 6 | 3 | 6.12 |
| Generalized Gini | 6 | 1 | 1 | 4 | 4 | 1 | 1 | 1 | 1 | 5 | 2 | 1 | 1 | 1 | 4 | 1 | 2.19 |
| JensenShannon | 12 | 12 | 12 | 1 | 12 | 4 | 11 | 5 | 2 | 4 | 12 | 12 | 12 | 11 | 11 | 10 | 8.94 |
| PearsonChi2 | 3 | 10 | 3 | 7 | 3 | 7 | 3 | 4 | 5 | 5 | 3 | 5 | 2 | 7 | 7 | 8 | 5.12 |
| VicisSymmetric1 | 10 | 9 | 11 | 5 | 2 | 11 | 3 | 8 | 9 | 1 | 11 | 11 | 11 | 12 | 3 | 10 | 7.94 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Euclidean | 9 | 1 | 10 | 9 | 2 | 2 | 8 | 10 | 2 | 11 | 8 | 7 | 5 | 8 | 10 | 6 | 6.75 |
| Manhattan | 3 | 1 | 9 | 7 | 4 | 6 | 7 | 10 | 8 | 10 | 6 | 5 | 8 | 3 | 8 | 3 | 6.12 |
| Minkowski | 8 | 1 | 8 | 11 | 8 | 6 | 10 | 9 | 2 | 12 | 9 | 9 | 5 | 10 | 11 | 4 | 7.69 |
| Cosine | 1 | 7 | 7 | 10 | 1 | 8 | 11 | 10 | 2 | 9 | 10 | 3 | 10 | 4 | 9 | 1 | 6.44 |
| Hassanat | 1 | 5 | 2 | 2 | 6 | 4 | 8 | 4 | 11 | 2 | 1 | 8 | 7 | 4 | 2 | 9 | 4.75 |
| Canberra | 7 | 8 | 6 | 1 | 12 | 4 | 6 | 3 | 11 | 2 | 7 | 11 | 9 | 6 | 1 | 11 | 6.56 |
| Hellinger | 5 | 10 | 4 | 3 | 5 | 10 | 2 | 6 | 2 | 7 | 3 | 3 | 4 | 9 | 5 | 7 | 5.31 |
| Gini | 11 | 5 | 4 | 8 | 6 | 12 | 5 | 8 | 8 | 8 | 5 | 2 | 3 | 2 | 7 | 4 | 6.12 |
| Generalized Gini | 5 | 1 | 1 | 4 | 9 | 1 | 1 | 2 | 1 | 5 | 2 | 1 | 1 | 1 | 4 | 2 | 2.56 |
| JensenShannon | 12 | 12 | 12 | 12 | 11 | 2 | 12 | 1 | 2 | 4 | 12 | 12 | 12 | 11 | 12 | 12 | 9.44 |
| PearsonChi2 | 4 | 10 | 3 | 6 | 3 | 10 | 3 | 7 | 2 | 5 | 3 | 6 | 2 | 7 | 6 | 8 | 5.31 |
| VicisSymmetric1 | 10 | 9 | 11 | 4 | 9 | 9 | 3 | 5 | 8 | 1 | 11 | 10 | 11 | 12 | 3 | 10 | 7.88 |
B.1.3 With noise at level 10%
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 2 | 0.885 | 0.880 | 2 | 1.000 | 1.000 | 2 | 0.687 | 0.652 | 1 | 0.565 | 0.571 | 8 | 0.850 | 0.875 | 8 | 0.747 | 0.765 | 9 | 0.902 | 0.902 | 1 | 0.680 | 0.665 |
| Manhattan | 2 | 0.900 | 0.900 | 1 | 0.999 | 0.999 | 2 | 0.701 | 0.665 | 4 | 0.565 | 0.594 | 9 | 0.815 | 0.875 | 10 | 0.743 | 0.761 | 8 | 0.911 | 0.909 | 1 | 0.678 | 0.663 |
| Minkowski | 2 | 0.888 | 0.886 | 2 | 1.000 | 1.000 | 4 | 0.686 | 0.663 | 1 | 0.549 | 0.558 | 8 | 0.843 | 0.869 | 8 | 0.764 | 0.775 | 9 | 0.895 | 0.893 | 1 | 0.685 | 0.671 |
| Cosine | 2 | 0.906 | 0.897 | 1 | 0.999 | 0.999 | 8 | 0.696 | 0.692 | 6 | 0.540 | 0.564 | 1 | 0.916 | 0.907 | 9 | 0.736 | 0.758 | 10 | 0.882 | 0.877 | 1 | 0.671 | 0.668 |
| Hassanat | 3 | 0.902 | 0.903 | 7 | 0.999 | 0.999 | 9 | 0.712 | 0.713 | 5 | 0.657 | 0.701 | 8 | 0.820 | 0.872 | 4 | 0.738 | 0.764 | 9 | 0.910 | 0.909 | 1 | 0.680 | 0.668 |
| Canberra | 6 | 0.895 | 0.891 | 9 | 0.994 | 0.993 | 5 | 0.696 | 0.694 | 7 | 0.682 | 0.721 | 10 | 0.823 | 0.829 | 4 | 0.742 | 0.764 | 9 | 0.918 | 0.914 | 1 | 0.674 | 0.711 |
| Hellinger | 2 | 0.902 | 0.902 | 1 | 0.944 | 0.944 | 9 | 0.712 | 0.704 | 5 | 0.631 | 0.671 | 9 | 0.850 | 0.872 | 5 | 0.737 | 0.752 | 5 | 0.928 | 0.924 | 1 | 0.699 | 0.694 |
| Gini | 2 | 0.849 | 0.840 | 2 | 0.999 | 0.999 | 7 | 0.711 | 0.702 | 5 | 0.599 | 0.580 | 9 | 0.845 | 0.873 | 3 | 0.704 | 0.735 | 4 | 0.918 | 0.916 | 1 | 0.680 | 0.675 |
| Generalized Gini | 3 4.91 | 0.886 | 0.883 | 2 1.59 | 1.000 | 1.000 | 8 2.94 | 0.712 | 0.712 | 4 3.22 | 0.658 | 0.653 | 3 4.18 | 0.857 | 0.850 | 6 1.21 | 0.734 | 0.758 | 8 3.56 | 0.929 | 0.927 | 5 4.49 | 0.698 | 0.710 |
| JensenShannon | 2 | 0.643 | 0.442 | 4 | 0.743 | 0.537 | 8 | 0.618 | 0.594 | 2 | 0.587 | 0.261 | 2 | 0.722 | 0.866 | 7 | 0.746 | 0.765 | 9 | 0.884 | 0.874 | 2 | 0.686 | 0.719 |
| PearsonChi2 | 2 | 0.898 | 0.897 | 1 | 0.945 | 0.945 | 8 | 0.724 | 0.712 | 3 | 0.632 | 0.608 | 8 | 0.855 | 0.875 | 4 | 0.741 | 0.751 | 7 | 0.924 | 0.920 | 1 | 0.692 | 0.686 |
| VicisSymmetric1 | 3 | 0.843 | 0.835 | 2 | 0.988 | 0.988 | 4 | 0.642 | 0.633 | 3 | 0.615 | 0.651 | 9 | 0.854 | 0.856 | 6 | 0.732 | 0.751 | 6 | 0.923 | 0.920 | 10 | 0.676 | 0.711 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Euclidean | 7 | 0.967 | 0.967 | 1 | 0.706 | 0.707 | 10 | 0.933 | 0.933 | 1 | 0.848 | 0.846 | 6 | 0.637 | 0.648 | 1 | 0.868 | 0.865 | 10 | 0.707 | 0.699 | 1 | 0.703 | 0.678 |
| Manhattan | 3 | 0.967 | 0.967 | 1 | 0.770 | 0.769 | 10 | 0.943 | 0.942 | 1 | 0.851 | 0.846 | 9 | 0.647 | 0.655 | 1 | 0.902 | 0.899 | 8 | 0.744 | 0.726 | 1 | 0.678 | 0.654 |
| Minkowski | 7 | 0.967 | 0.967 | 1 | 0.677 | 0.679 | 10 | 0.933 | 0.933 | 1 | 0.838 | 0.832 | 5 3 | 0.621 | 0.633 | 1 | 0.851 | 0.847 | 10 | 0.707 | 0.699 | 1 | 0.725 | 0.696 |
| Cosine | 7 | 0.965 | 0.960 | 1 | 0.835 | 0.831 | 7 | 0.924 | 0.924 | 1 | 0.838 | 0.832 | 4 | 0.622 | 0.634 | 1 | 0.884 | 0.884 | 9 | 0.716 | 0.715 | 1 | 0.720 | 0.710 |
| Hassanat | 3 | 0.960 | 0.960 | 7 | 0.960 | 0.955 | 6 | 0.955 | 0.954 | 1 | 0.840 | 0.837 | 7 | 0.645 | 0.653 | 7 | 0.884 | 0.882 | 8 | 0.844 | 0.841 | 4 | 0.685 | 0.645 |
| Canberra | 7 | 0.960 | 0.960 | 7 | 0.958 | 0.955 | 9 | 0.939 | 0.938 | 10 | 0.793 | 0.658 | 9 | 0.623 | 0.620 | 10 | 0.869 | 0.869 | 10 | 0.863 | 0.862 | 1 | 0.514 | 0.509 |
| Hellinger | 5 | 0.960 | 0.960 | 4 | 0.918 | 0.915 | 4 | 0.946 | 0.945 | 1 | 0.838 | 0.835 | 8 | 0.639 | 0.647 | 1 | 0.861 | 0.861 | 5 | 0.764 | 0.763 | 3 | 0.664 | 0.668 |
| Gini | 3 | 0.960 | 0.960 | 6 | 0.898 | 0.893 | 3 | 0.944 | 0.943 | 1 | 0.838 | 0.835 | 8 | 0.655 | 0.660 | 3 | 0.924 | 0.923 | 4 | 0.746 | 0.742 | 4 | 0.736 | 0.696 |
| Generalized Gini | 3 1.74 | 0.967 | 0.967 | 3 4.93 | 0.914 | 0.910 | 2 1.11 | 0.950 | 0.949 | 1 1.47 | 0.873 | 0.870 | 9 2.92 | 0.666 | 0.666 | 1 4.45 | 0.933 | 0.931 | 4 3.23 | 0.791 | 0.773 | 2 1.69 | 0.736 | 0.672 |
| JensenShannon | 1 | 0.953 | 0.953 | 3 | 0.950 | 0.949 | 4 | 0.717 | 0.649 | 4 | 0.285 | 0.534 | 4 | 0.519 | 0.588 | 3 | 0.840 | 0.834 | 9 | 0.708 | 0.649 | 1 | 0.553 | 0.379 |
| PearsonChi2 | 3 | 0.960 | 0.960 | 4 | 0.918 | 0.915 | 4 | 0.946 | 0.945 | 1 | 0.838 | 0.832 | 9 | 0.656 | 0.665 | 4 | 0.864 | 0.864 | 7 | 0.755 | 0.740 | 4 | 0.668 | 0.654 |
| VicisSymmetric1 | 9 | 0.980 | 0.980 | 9 | 0.951 | 0.944 | 6 | 0.887 | 0.886 | 9 | 0.772 | 0.735 | 7 | 0.597 | 0.605 | 3 | 0.834 | 0.829 | 3 | 0.829 | 0.828 | 1 | 0.574 | 0.566 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Euclidean | 9 | 1 | 9 | 9 | 5 | 2 | 9 | 6 | 2 | 11 | 8 | 3 | 7 | 7 | 11 | 5 | 6.5 |
| Manhattan | 4 | 4 | 6 | 9 | 11 | 4 | 7 | 9 | 2 | 10 | 6 | 2 | 4 | 3 | 8 | 7 | 6 |
| Minkowski | 7 | 1 | 10 | 11 | 8 | 1 | 10 | 5 | 2 | 12 | 8 | 5 | 10 | 10 | 11 | 3 | 7.12 |
| Cosine | 1 | 4 | 7 | 12 | 1 | 9 | 12 | 12 | 6 | 9 | 10 | 5 | 9 | 4 | 9 | 4 | 7.12 |
| Hassanat | 2 | 4 | 2 | 3 | 10 | 7 | 8 | 6 | 7 | 1 | 1 | 4 | 5 | 4 | 2 | 6 | 4.5 |
| Canberra | 6 | 8 | 7 | 1 | 9 | 5 | 5 | 11 | 7 | 2 | 7 | 10 | 8 | 6 | 1 | 12 | 6.56 |
| Hellinger | 2 | 11 | 2 | 5 | 5 | 8 | 2 | 1 | 7 | 5 | 3 | 5 | 6 | 9 | 5 | 9 | 5.31 |
| Gini | 10 | 4 | 5 | 7 | 7 | 12 | 5 | 6 | 7 | 8 | 5 | 5 | 3 | 2 | 7 | 1 | 5.88 |
| Generalized Gini | 8 | 1 | 2 | 2 | 2 | 10 | 1 | 2 | 2 | 7 | 2 | 1 | 1 | 1 | 4 | 1 | 2.94 |
| JensenShannon | 12 | 12 | 12 | 8 | 12 | 3 | 11 | 4 | 12 | 4 | 12 | 12 | 12 | 11 | 10 | 11 | 9.88 |
| PearsonChi2 | 5 | 10 | 1 | 4 | 3 | 6 | 3 | 3 | 7 | 5 | 3 | 5 | 2 | 8 | 6 | 8 | 4.94 |
| VicisSymmetric1 | 11 | 9 | 11 | 6 | 4 | 11 | 4 | 10 | 1 | 3 | 11 | 11 | 11 | 12 | 3 | 10 | 8 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Euclidean | 9 | 1 | 10 | 9 | 2 | 2 | 9 | 11 | 2 | 11 | 8 | 2 | 6 | 7 | 10 | 4 | 6.44 |
| Manhattan | 3 | 4 | 8 | 7 | 2 | 6 | 7 | 12 | 2 | 10 | 6 | 2 | 4 | 3 | 8 | 7 | 5.69 |
| Minkowski | 7 | 1 | 9 | 11 | 8 | 1 | 10 | 8 | 2 | 12 | 8 | 7 | 9 | 10 | 10 | 2 | 7.19 |
| Cosine | 4 | 4 | 7 | 10 | 1 | 7 | 11 | 9 | 6 | 9 | 10 | 7 | 8 | 4 | 9 | 1 | 6.69 |
| Hassanat | 1 | 4 | 1 | 2 | 6 | 4 | 7 | 9 | 6 | 1 | 1 | 4 | 5 | 5 | 2 | 9 | 4.19 |
| Canberra | 6 | 8 | 6 | 1 | 12 | 4 | 6 | 2 | 6 | 1 | 7 | 11 | 10 | 6 | 1 | 11 | 6.12 |
| Hellinger | 2 | 11 | 4 | 3 | 6 | 9 | 2 | 5 | 6 | 5 | 3 | 5 | 7 | 9 | 5 | 6 | 5.5 |
| Gini | 10 | 4 | 5 | 8 | 5 | 12 | 5 | 7 | 6 | 8 | 5 | 5 | 3 | 2 | 6 | 2 | 5.81 |
| Generalized Gini | 8 | 1 | 2 | 4 | 11 | 7 | 1 | 4 | 2 | 7 | 2 | 1 | 1 | 1 | 4 | 5 | 3.81 |
| JensenShannon | 12 | 12 | 12 | 12 | 9 | 2 | 12 | 1 | 12 | 3 | 12 | 12 | 12 | 11 | 12 | 12 | 9.88 |
| PearsonChi2 | 4 | 10 | 2 | 6 | 2 | 10 | 3 | 6 | 6 | 5 | 3 | 7 | 2 | 8 | 7 | 7 | 5.5 |
| VicisSymmetric1 | 11 | 9 | 11 | 5 | 10 | 10 | 3 | 2 | 1 | 4 | 11 | 10 | 11 | 12 | 3 | 10 | 7.69 |
B.2 KMeans experiments
B.2.1 Without noise
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 4 | 1 | 1 | 1 | 1 | 5 | 6 | 3 | 4 | 6 | 3 | 1 | 8 | 6 | 2 | 4 | 3.50 |
| Gini | 7 | 8 | 4 | 5 | 4 | 7 | 12 | 6 | 5 | 8 | 4 | 2 | 12 | 7 | 3 | 8 | 6.38 |
| Euclidean | 10 | 4 | 5 | 8 | 5 | 11 | 3 | 10 | 8 | 11 | 8 | 7 | 10 | 10 | 10 | 10 | 8.12 |
| Manhattan | 8 | 3 | 5 | 10 | 3 | 9 | 1 | 10 | 6 | 9 | 9 | 5 | 9 | 11 | 10 | 6 | 7.12 |
| Minkowski | 9 | 2 | 5 | 7 | 6 | 10 | 5 | 10 | 7 | 10 | 7 | 6 | 11 | 9 | 10 | 11 | 7.81 |
| Cosine | 6 | 6 | 10 | 3 | 2 | 8 | 2 | 8 | 12 | 7 | 5 | 8 | 6 | 12 | 5 | 7 | 6.69 |
| Canberra | 2 | 7 | 3 | 2 | 9 | 1 | 11 | 5 | 11 | 2 | 2 | 10 | 4 | 8 | 9 | 5 | 5.69 |
| Hellinger | 1 | 5 | 8 | 9 | 8 | 6 | 9 | 1 | 2 | 4 | 11 | 12 | 1 | 2 | 7 | 2 | 5.5 |
| Jensen-Shannon | 11 | 12 | 9 | 12 | 10 | 12 | 10 | 1 | 1 | 12 | 12 | 9 | 2 | 3 | 4 | 12 | 8.25 |
| PearsonChi2 | 11 | 11 | 12 | 6 | 11 | 3 | 4 | 7 | 3 | 1 | 6 | 3 | 5 | 4 | 6 | 1 | 5.88 |
| VicisSymmetric1 | 3 | 10 | 11 | 11 | 12 | 2 | 8 | 9 | 9 | 5 | 10 | 11 | 3 | 1 | 8 | 3 | 7.25 |
| Hassanat | 5 | 9 | 2 | 4 | 7 | 4 | 7 | 4 | 10 | 2 | 1 | 4 | 7 | 5 | 1 | 9 | 5.06 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 4 | 1 | 2 | 2 | 2 | 9 | 7 | 7 | 4 | 6 | 2 | 1 | 6 | 6 | 4 | 4 | 4.19 |
| Gini | 9 | 8 | 3 | 3 | 3 | 7 | 7 | 6 | 4 | 9 | 4 | 3 | 8 | 7 | 5 | 6 | 5.75 |
| Euclidean | 12 | 4 | 5 | 5 | 4 | 10 | 10 | 10 | 7 | 11 | 8 | 8 | 11 | 9 | 9 | 10 | 8.31 |
| Manhattan | 10 | 2 | 5 | 8 | 1 | 8 | 12 | 10 | 4 | 10 | 9 | 5 | 9 | 10 | 9 | 8 | 7.5 |
| Minkowski | 11 | 2 | 5 | 7 | 6 | 11 | 6 | 10 | 7 | 11 | 7 | 7 | 12 | 8 | 9 | 10 | 8.06 |
| Cosine | 6 | 5 | 11 | 4 | 5 | 6 | 2 | 4 | 12 | 8 | 5 | 9 | 4 | 12 | 2 | 7 | 6.38 |
| Canberra | 3 | 6 | 4 | 9 | 10 | 3 | 11 | 2 | 10 | 3 | 1 | 11 | 2 | 10 | 12 | 5 | 6.38 |
| Hellinger | 2 | 7 | 12 | 10 | 9 | 4 | 4 | 9 | 2 | 4 | 11 | 12 | 3 | 4 | 7 | 1 | 6.31 |
| Jensen-Shannon | 7 | 9 | 8 | 12 | 12 | 12 | 5 | 8 | 1 | 7 | 12 | 4 | 1 | 1 | 3 | 12 | 7.12 |
| PearsonChi2 | 7 | 12 | 10 | 6 | 11 | 2 | 1 | 3 | 3 | 1 | 6 | 2 | 5 | 3 | 6 | 2 | 5 |
| VicisSymmetric1 | 1 | 11 | 9 | 11 | 8 | 5 | 9 | 1 | 11 | 5 | 10 | 10 | 7 | 2 | 8 | 3 | 6.94 |
| Hassanat | 5 | 10 | 1 | 1 | 7 | 1 | 3 | 5 | 9 | 2 | 2 | 6 | 10 | 5 | 1 | 8 | 4.75 |
B.2.2 With noise at level 5%
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini_generalized | 1.1 | 0.5391 | 0.4359 | 4.92 | 0.6306 | 0.6334 | 2.40 | 0.5501 | 0.5470 | 2.32 | 0.5959 | 0.3566 | 2.44 | 0.6283 | 0.4736 | 1.82 | 0.5561 | 0.3499 | 2.30 | 0.2688 | 0.1364 | 1.58 | 0.7066 | 0.0552 |
| Gini | - | 0.3996 | 0.3447 | - | 0.5779 | 0.5773 | - | 0.5140 | 0.5140 | - | 0.5373 | 0.3633 | - | 0.5660 | 0.4336 | - | 0.5344 | 0.3432 | - | 0.1457 | 0.1341 | - | 0.6802 | 0.1054 |
| Euclidean | - | 0.3344 | 0.2906 | - | 0.6070 | 0.6115 | - | 0.5120 | 0.5120 | - | 0.5136 | 0.2574 | - | 0.5975 | 0.4576 | - | 0.5479 | 0.3335 | - | 0.5428 | 0.1250 | - | 0.7150 | 0.0155 |
| Manhattan | - | 0.3792 | 0.3276 | - | 0.6071 | 0.6115 | - | 0.5110 | 0.5110 | - | 0.4994 | 0.2540 | - | 0.6476 | 0.4912 | - | 0.5332 | 0.3335 | - | 0.5508 | 0.0977 | - | 0.7150 | 0.0155 |
| Minkowski | 3.0 | 0.3309 | 0.2878 | 3.0 | 0.6072 | 0.6115 | 3.0 | 0.5120 | 0.5120 | 3.0 | 0.5045 | 0.2607 | 3.0 | 0.6127 | 0.4752 | 3.0 | 0.5515 | 0.3204 | 3.0 | 0.5340 | 0.1409 | 3.0 | 0.7150 | 0.0155 |
| Cosine | - | 0.4291 | 0.3645 | - | 0.5976 | 0.6006 | - | 0.4885 | 0.4900 | - | 0.5169 | 0.2708 | - | 0.6354 | 0.4880 | - | 0.5452 | 0.3627 | - | 0.5901 | 0.4909 | - | 0.6305 | 0.1640 |
| Canberra | - | 0.5411 | 0.4386 | - | 0.5980 | 0.6021 | - | 0.5439 | 0.5300 | - | 0.4764 | 0.1987 | - | 0.6047 | 0.4592 | - | 0.5521 | 0.3920 | - | 0.2220 | 0.1068 | - | 0.6860 | 0.1745 |
| Hellinger | - | 0.7727 | 0.6640 | - | 0.5949 | 0.5860 | - | 0.4916 | 0.4900 | - | 0.4628 | 0.2242 | - | 0.4992 | 0.3440 | - | 0.5426 | 0.4308 | - | 0.2941 | 0.2136 | - | 0.7150 | 0.0432 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.2500 | 0.5000 | - | 0.1495 | 0.0132 | - | 0.5386 | 0.1872 | - | 0.5328 | 0.0918 | - | 0.2626 | 0.1409 | - | 0.7150 | 0.0518 |
| PearsonChi2 | - | 0.5953 | 0.5153 | - | 0.5111 | 0.5109 | - | 0.5030 | 0.5030 | - | 0.4480 | 0.2309 | - | 0.4977 | 0.4224 | - | 0.4466 | 0.2981 | - | 0.5010 | 0.4909 | - | 0.6051 | 0.2381 |
| VicisSymmetric1 | - | 0.5344 | 0.4727 | - | 0.5036 | 0.5014 | - | 0.4917 | 0.5070 | - | 0.4717 | 0.2345 | - | 0.4313 | 0.3744 | - | 0.5110 | 0.3162 | - | 0.3415 | 0.2864 | - | 0.5583 | 0.2870 |
| Hassanat | - | 0.4426 | 0.3760 | - | 0.5572 | 0.5539 | - | 0.5619 | 0.5610 | - | 0.5826 | 0.3333 | - | 0.5862 | 0.4496 | - | 0.5901 | 0.5520 | - | 0.3647 | 0.3091 | - | 0.6915 | 0.1486 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini_generalized | 2.68 | 0.0939 | 0.1067 | 4.56 | 0.2085 | 0.2021 | 2.24 | 0.9283 | 0.9245 | 2.04 | 0.5750 | 0.5722 | 2.30 | 0.3871 | 0.2850 | 4.76 | 0.3050 | 0.2872 | 1.40 | 0.7677 | 0.6000 | 1.50 | 0.0918 | 0.0514 |
| Gini | - | 0.0761 | 0.1000 | - | 0.1207 | 0.1010 | - | 0.9107 | 0.9034 | - | 0.5460 | 0.5431 | - | 0.3542 | 0.2684 | - | 0.2654 | 0.2576 | - | 0.7086 | 0.6145 | - | 0.0774 | 0.0421 |
| Euclidean | - | 0.0728 | 0.0933 | - | 0.0610 | 0.0449 | - | 0.8771 | 0.8489 | - | 0.5448 | 0.5382 | - | 0.3563 | 0.2337 | - | 0.2105 | 0.2292 | - | 0.4877 | 0.5594 | - | 0.0358 | 0.0188 |
| Manhattan | - | 0.0761 | 0.1000 | - | 0.0554 | 0.0394 | - | 0.8735 | 0.8437 | - | 0.5621 | 0.5575 | - | 0.3669 | 0.2632 | - | 0.2055 | 0.2221 | - | 0.4877 | 0.5594 | - | 0.0455 | 0.0235 |
| Minkowski | 3.0 | 0.0690 | 0.0867 | 3.0 | 0.0610 | 0.0449 | 3.0 | 0.8793 | 0.8524 | 3.0 | 0.5331 | 0.5240 | 3.0 | 0.3552 | 0.2325 | 3.0 | 0.2098 | 0.2292 | 3.0 | 0.4877 | 0.5594 | 3.0 | 0.0336 | 0.0188 |
| Cosine | - | 0.0211 | 0.0200 | - | 0.1536 | 0.1235 | - | 0.9131 | 0.9034 | - | 0.5460 | 0.5431 | - | 0.3990 | 0.3081 | - | 0.1968 | 0.2162 | - | 0.6776 | 0.6725 | - | 0.0554 | 0.0467 |
| Canberra | - | 0.0539 | 0.0667 | - | 0.3879 | 0.3483 | - | 0.9223 | 0.9209 | - | 0.6073 | 0.6062 | - | 0.4330 | 0.3402 | - | 0.2386 | 0.2304 | - | 0.5703 | 0.5913 | - | 0.0806 | 0.0470 |
| Hellinger | - | 0.1529 | 0.1533 | - | 0.4685 | 0.4213 | - | 0.7039 | 0.6893 | - | 0.6826 | 0.5002 | - | 0.4312 | 0.3348 | - | 0.4326 | 0.3630 | - | 0.6697 | 0.5710 | - | 0.2237 | 0.1401 |
| Jensen-Shannon | - | 0.4233 | 0.4467 | - | 0.0490 | 0.1183 | - | 0.1380 | 0.3691 | - | 0.2849 | 0.5337 | - | 0.3487 | 0.5905 | - | 0.3686 | 0.4350 | - | 0.3864 | 0.5623 | - | 0.0000 | 0.0000 |
| PearsonChi2 | - | 0.2017 | 0.1867 | - | 0.3128 | 0.3830 | - | 0.8853 | 0.8647 | - | 0.4072 | 0.4086 | - | 0.4221 | 0.3261 | - | 0.2652 | 0.2351 | - | 0.7337 | 0.5855 | - | 0.2041 | 0.1728 |
| VicisSymmetric1 | - | 0.0667 | 0.0267 | - | 0.4796 | 0.4622 | - | 0.5351 | 0.5009 | - | 0.5330 | 0.5283 | - | 0.4063 | 0.2567 | - | 0.3746 | 0.3275 | - | 0.3966 | 0.4986 | - | 0.2229 | 0.1353 |
| Hassanat | - | 0.0637 | 0.0800 | - | 0.3990 | 0.3541 | - | 0.9245 | 0.9157 | - | 0.5578 | 0.5527 | - | 0.3962 | 0.2581 | - | 0.3502 | 0.2943 | - | 0.7850 | 0.7710 | - | 0.0351 | 0.0234 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 4 | 1 | 2 | 1 | 3 | 2 | 9 | 6 | 4 | 6 | 1 | 3 | 7 | 5 | 2 | 4 | 3.75 |
| Gini | 8 | 8 | 4 | 3 | 8 | 8 | 12 | 9 | 5 | 8 | 5 | 6 | 11 | 6 | 4 | 6 | 6.94 |
| Euclidean | 10 | 4 | 5 | 5 | 6 | 5 | 3 | 1 | 7 | 9 | 8 | 8 | 9 | 9 | 8 | 9 | 6.62 |
| Manhattan | 9 | 3 | 7 | 7 | 1 | 9 | 2 | 1 | 5 | 11 | 9 | 4 | 8 | 11 | 8 | 8 | 6.44 |
| Minkowski | 11 | 2 | 5 | 6 | 4 | 4 | 4 | 1 | 8 | 9 | 7 | 9 | 10 | 10 | 8 | 11 | 6.81 |
| Cosine | 7 | 6 | 11 | 4 | 2 | 6 | 1 | 10 | 12 | 7 | 4 | 6 | 5 | 12 | 5 | 7 | 6.56 |
| Canberra | 3 | 5 | 3 | 8 | 5 | 3 | 11 | 8 | 11 | 4 | 3 | 2 | 1 | 8 | 7 | 5 | 5.44 |
| Hellinger | 1 | 7 | 10 | 10 | 10 | 7 | 8 | 1 | 3 | 2 | 10 | 1 | 2 | 1 | 6 | 1 | 5 |
| JensenShannon | 12 | 12 | 12 | 12 | 9 | 10 | 10 | 1 | 1 | 12 | 12 | 12 | 12 | 3 | 12 | 12 | 9.62 |
| PearsonChi2 | 2 | 10 | 8 | 11 | 11 | 12 | 5 | 11 | 2 | 5 | 6 | 11 | 3 | 7 | 3 | 3 | 6.88 |
| VicisSymmetric1 | 5 | 11 | 9 | 9 | 12 | 11 | 7 | 12 | 9 | 1 | 11 | 10 | 4 | 2 | 11 | 2 | 7.88 |
| Hassanat | 6 | 9 | 1 | 2 | 7 | 1 | 6 | 7 | 10 | 3 | 2 | 5 | 6 | 4 | 1 | 10 | 5 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 5 | 1 | 2 | 2 | 4 | 5 | 8 | 7 | 4 | 6 | 1 | 2 | 6 | 5 | 4 | 4 | 4.12 |
| Gini | 9 | 8 | 4 | 1 | 8 | 6 | 9 | 6 | 5 | 9 | 4 | 5 | 7 | 6 | 3 | 7 | 6.06 |
| Euclidean | 11 | 2 | 5 | 6 | 6 | 7 | 10 | 10 | 7 | 10 | 8 | 7 | 11 | 9 | 9 | 10 | 8 |
| Manhattan | 10 | 2 | 7 | 7 | 1 | 7 | 12 | 10 | 5 | 12 | 9 | 3 | 8 | 11 | 9 | 8 | 7.56 |
| Minkowski | 12 | 2 | 5 | 5 | 3 | 9 | 6 | 10 | 8 | 10 | 7 | 10 | 12 | 9 | 9 | 10 | 7.94 |
| Cosine | 7 | 6 | 11 | 4 | 2 | 4 | 1 | 4 | 12 | 7 | 4 | 5 | 5 | 12 | 2 | 6 | 5.75 |
| Canberra | 4 | 5 | 3 | 11 | 5 | 3 | 11 | 3 | 10 | 5 | 2 | 1 | 2 | 8 | 5 | 5 | 5.19 |
| Hellinger | 1 | 7 | 11 | 10 | 11 | 2 | 5 | 9 | 3 | 2 | 10 | 11 | 3 | 2 | 7 | 2 | 6 |
| Jensen-Shannon | 8 | 9 | 10 | 12 | 12 | 12 | 6 | 8 | 1 | 8 | 12 | 8 | 1 | 1 | 8 | 12 | 8 |
| PearsonChi2 | 2 | 11 | 9 | 9 | 9 | 11 | 1 | 2 | 2 | 3 | 6 | 12 | 4 | 7 | 6 | 1 | 5.94 |
| VicisSymmetric1 | 3 | 12 | 8 | 8 | 10 | 10 | 4 | 1 | 11 | 1 | 11 | 9 | 10 | 3 | 12 | 3 | 7.25 |
| Hassanat | 6 | 10 | 1 | 3 | 7 | 1 | 3 | 5 | 9 | 4 | 3 | 4 | 9 | 4 | 1 | 9 | 4.94 |
B.2.3 With noise at level 10%
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini_generalized | 3.38 | 0.0946 | 0.1067 | 4.78 | 0.2116 | 0.2022 | 2.24 | 0.9283 | 0.9245 | 2.94 | 0.5864 | 0.5815 | 2.14 | 0.3871 | 0.2863 | 4.44 | 0.3080 | 0.2872 | 1.34 | 0.7672 | 0.5986 | 2.38 | 0.1010 | 0.0560 |
| Gini | - | 0.0768 | 0.1000 | - | 0.1207 | 0.1010 | - | 0.9107 | 0.9034 | - | 0.5586 | 0.5527 | - | 0.3614 | 0.2722 | - | 0.2651 | 0.2576 | - | 0.7101 | 0.6159 | - | 0.0356 | 0.0186 |
| Euclidean | - | 0.0806 | 0.1000 | - | 0.0610 | 0.0449 | - | 0.8771 | 0.8489 | - | 0.5536 | 0.5479 | - | 0.3571 | 0.2350 | - | 0.2097 | 0.2281 | - | 0.4877 | 0.5594 | - | 0.0435 | 0.0235 |
| Manhattan | - | 0.0761 | 0.1000 | - | 0.0554 | 0.0394 | - | 0.8735 | 0.8437 | - | 0.5488 | 0.5432 | - | 0.3640 | 0.2581 | - | 0.2055 | 0.2221 | - | 0.4877 | 0.5594 | - | 0.0502 | 0.0281 |
| Minkowski | 3.0 | 0.0751 | 0.0933 | 3.0 | 0.0610 | 0.0449 | 3.0 | 0.8793 | 0.8524 | 3.0 | 0.5742 | 0.5672 | 3.0 | 0.3551 | 0.2325 | 3.0 | 0.2105 | 0.2304 | 3.0 | 0.4877 | 0.5594 | 3.0 | 0.0435 | 0.0235 |
| Cosine | - | 0.0216 | 0.0200 | - | 0.1514 | 0.1235 | - | 0.9131 | 0.9034 | - | 0.5523 | 0.5479 | - | 0.3989 | 0.3056 | - | 0.1968 | 0.2162 | - | 0.6776 | 0.6725 | - | 0.0626 | 0.0467 |
| Canberra | - | 0.0303 | 0.0333 | - | 0.3879 | 0.3484 | - | 0.9223 | 0.9209 | - | 0.5474 | 0.5483 | - | 0.4397 | 0.3543 | - | 0.2333 | 0.2257 | - | 0.5599 | 0.5696 | - | 0.1254 | 0.0699 |
| Hellinger | - | 0.3318 | 0.3600 | - | 0.5159 | 0.4602 | - | 0.7039 | 0.6893 | - | 0.5904 | 0.4710 | - | 0.4265 | 0.3416 | - | 0.4041 | 0.3617 | - | 0.6697 | 0.5710 | - | 0.2508 | 0.1355 |
| Jensen-Shannon | - | 0.4514 | 0.4800 | - | 0.0525 | 0.1295 | - | 0.1388 | 0.3726 | - | 0.2849 | 0.5337 | - | 0.4401 | 0.5661 | - | 0.3701 | 0.4373 | - | 0.4003 | 0.5667 | - | 0.0000 | 0.0000 |
| PearsonChi2 | - | 0.2908 | 0.2667 | - | 0.2757 | 0.2408 | - | 0.8904 | 0.8683 | - | 0.4979 | 0.4906 | - | 0.5424 | 0.4004 | - | 0.3915 | 0.2918 | - | 0.7548 | 0.5826 | - | 0.2871 | 0.1543 |
| VicisSymmetric1 | - | 0.0000 | 0.0000 | - | 0.2393 | 0.2359 | - | 0.5108 | 0.4905 | - | 0.4841 | 0.4764 | - | 0.4278 | 0.3028 | - | 0.3789 | 0.3238 | - | 0.3890 | 0.4841 | - | 0.2037 | 0.1351 |
| Hassanat | - | 0.0457 | 0.0533 | - | 0.3931 | 0.3484 | - | 0.9237 | 0.9174 | - | 0.5488 | 0.5432 | - | 0.3917 | 0.2568 | - | 0.3494 | 0.2919 | - | 0.7838 | 0.7696 | - | 0.0629 | 0.0374 |
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini_generalized | 1.1 | 0.4815 | 0.4045 | 4.98 | 0.6305 | 0.6334 | 2.38 | 0.5510 | 0.5480 | 2.32 | 0.5966 | 0.3532 | 2.16 | 0.6248 | 0.4608 | 1.74 | 0.5615 | 0.3435 | 2.30 | 0.2688 | 0.1364 | 1.56 | 0.7066 | 0.0552 |
| Gini | - | 0.3924 | 0.3389 | - | 0.5787 | 0.5780 | - | 0.5140 | 0.5140 | - | 0.5389 | 0.3601 | - | 0.5608 | 0.4336 | - | 0.5451 | 0.3433 | - | 0.1457 | 0.1341 | - | 0.6802 | 0.1054 |
| Euclidean | - | 0.3351 | 0.2905 | - | 0.6032 | 0.6079 | - | 0.5120 | 0.5120 | - | 0.5136 | 0.2574 | - | 0.6109 | 0.4688 | - | 0.5425 | 0.3302 | - | 0.5428 | 0.1250 | - | 0.7150 | 0.0155 |
| Manhattan | - | 0.3825 | 0.3304 | - | 0.6081 | 0.6123 | - | 0.5120 | 0.5120 | - | 0.5040 | 0.2540 | - | 0.6367 | 0.4768 | - | 0.5332 | 0.3335 | - | 0.5508 | 0.0977 | - | 0.7150 | 0.0155 |
| Minkowski | 3.0 | 0.3342 | 0.2906 | 3.0 | 0.6026 | 0.6072 | 3.0 | 0.5120 | 0.5120 | 3.0 | 0.5018 | 0.2607 | 3.0 | 0.6067 | 0.4864 | 3.0 | 0.5529 | 0.3237 | 3.0 | 0.5340 | 0.1409 | 3.0 | 0.7150 | 0.0155 |
| Cosine | - | 0.4329 | 0.3673 | - | 0.5976 | 0.6006 | - | 0.4885 | 0.4900 | - | 0.5169 | 0.2708 | - | 0.6290 | 0.4800 | - | 0.5452 | 0.3627 | - | 0.5901 | 0.4909 | - | 0.6305 | 0.1640 |
| Canberra | - | 0.4749 | 0.3959 | - | 0.5967 | 0.6013 | - | 0.6016 | 0.5900 | - | 0.6551 | 0.2808 | - | 0.5712 | 0.4352 | - | 0.5444 | 0.3854 | - | 0.2182 | 0.1023 | - | 0.6871 | 0.1797 |
| Hellinger | - | 0.6010 | 0.5212 | - | 0.5944 | 0.5860 | - | 0.4828 | 0.4840 | - | 0.6006 | 0.1911 | - | 0.4499 | 0.3632 | - | 0.5125 | 0.3472 | - | 0.3535 | 0.2909 | - | 0.7150 | 0.0432 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.2500 | 0.5000 | - | 0.0001 | 0.0066 | - | 0.6888 | 0.2576 | - | 0.4546 | 0.0754 | - | 0.2999 | 0.1477 | - | 0.7150 | 0.0535 |
| PearsonChi2 | - | 0.5366 | 0.4931 | - | 0.4841 | 0.4753 | - | 0.4909 | 0.4910 | - | 0.5051 | 0.2276 | - | 0.5122 | 0.3776 | - | 0.5686 | 0.4965 | - | 0.4801 | 0.4818 | - | 0.5753 | 0.3588 |
| VicisSymmetric1 | - | 0.4569 | 0.4017 | - | 0.5169 | 0.5131 | - | 0.4917 | 0.5070 | - | 0.4001 | 0.2474 | - | 0.4995 | 0.4064 | - | 0.5043 | 0.3392 | - | 0.3655 | 0.3318 | - | 0.6194 | 0.2139 |
| Hassanat | - | 0.4245 | 0.3618 | - | 0.5560 | 0.5525 | - | 0.5352 | 0.5350 | - | 0.6532 | 0.3668 | - | 0.6169 | 0.4784 | - | 0.5901 | 0.5520 | - | 0.3596 | 0.3000 | - | 0.6912 | 0.1520 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 3 | 1 | 2 | 4 | 4 | 3 | 10 | 6 | 4 | 6 | 1 | 2 | 8 | 6 | 2 | 5 | 4.19 |
| Gini | 8 | 8 | 4 | 5 | 9 | 6 | 12 | 9 | 6 | 8 | 5 | 4 | 10 | 7 | 4 | 11 | 7.25 |
| Euclidean | 10 | 3 | 5 | 7 | 6 | 8 | 3 | 1 | 5 | 9 | 8 | 5 | 11 | 10 | 8 | 9 | 6.75 |
| Manhattan | 9 | 2 | 5 | 9 | 2 | 9 | 2 | 1 | 7 | 11 | 9 | 7 | 9 | 11 | 8 | 8 | 6.81 |
| Minkowski | 11 | 4 | 5 | 10 | 7 | 4 | 4 | 1 | 8 | 9 | 7 | 3 | 12 | 9 | 8 | 9 | 6.94 |
| Cosine | 6 | 5 | 10 | 6 | 3 | 5 | 1 | 10 | 11 | 7 | 4 | 6 | 6 | 12 | 5 | 7 | 6.5 |
| Canberra | 4 | 6 | 1 | 1 | 8 | 7 | 11 | 8 | 10 | 3 | 3 | 9 | 3 | 8 | 7 | 4 | 5.81 |
| Hellinger | 1 | 7 | 11 | 3 | 12 | 10 | 8 | 1 | 2 | 1 | 10 | 1 | 5 | 1 | 6 | 2 | 5.06 |
| JensenShannon | 12 | 12 | 12 | 12 | 1 | 12 | 9 | 1 | 1 | 12 | 12 | 12 | 2 | 4 | 11 | 12 | 8.56 |
| PearsonChi2 | 2 | 11 | 9 | 8 | 10 | 2 | 5 | 12 | 3 | 4 | 6 | 10 | 1 | 2 | 3 | 1 | 5.56 |
| VicisSymmetric1 | 5 | 10 | 8 | 11 | 11 | 11 | 6 | 11 | 12 | 5 | 11 | 11 | 4 | 3 | 12 | 3 | 8.38 |
| Hassanat | 7 | 9 | 3 | 2 | 5 | 1 | 7 | 7 | 9 | 2 | 2 | 7 | 7 | 5 | 1 | 6 | 5 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 3 | 1 | 2 | 3 | 6 | 6 | 8 | 7 | 4 | 6 | 1 | 1 | 7 | 6 | 4 | 5 | 4.38 |
| Gini | 9 | 8 | 4 | 2 | 8 | 7 | 9 | 6 | 5 | 9 | 4 | 3 | 8 | 7 | 3 | 11 | 6.44 |
| Euclidean | 12 | 3 | 5 | 7 | 5 | 10 | 10 | 10 | 5 | 10 | 8 | 5 | 11 | 9 | 9 | 9 | 8 |
| Manhattan | 10 | 2 | 5 | 8 | 4 | 9 | 12 | 10 | 5 | 12 | 9 | 7 | 9 | 11 | 9 | 8 | 8.12 |
| Minkowski | 11 | 4 | 5 | 6 | 1 | 11 | 7 | 10 | 8 | 10 | 7 | 2 | 12 | 8 | 9 | 9 | 7.5 |
| Cosine | 6 | 6 | 11 | 5 | 2 | 4 | 1 | 4 | 11 | 8 | 4 | 5 | 5 | 12 | 2 | 6 | 5.75 |
| Canberra | 5 | 5 | 1 | 4 | 7 | 3 | 11 | 3 | 10 | 2 | 2 | 4 | 3 | 10 | 7 | 4 | 5.06 |
| Hellinger | 1 | 7 | 12 | 11 | 11 | 5 | 5 | 9 | 2 | 1 | 10 | 12 | 4 | 2 | 6 | 2 | 6.25 |
| JensenShannon | 8 | 9 | 9 | 12 | 12 | 12 | 6 | 8 | 1 | 7 | 12 | 9 | 1 | 1 | 8 | 12 | 7.94 |
| PearsonChi2 | 2 | 12 | 10 | 10 | 10 | 2 | 2 | 1 | 3 | 4 | 6 | 10 | 2 | 5 | 5 | 1 | 5.31 |
| VicisSymmetric1 | 4 | 11 | 8 | 9 | 9 | 8 | 3 | 2 | 12 | 5 | 11 | 11 | 6 | 3 | 12 | 3 | 7.31 |
| Hassanat | 7 | 10 | 3 | 1 | 3 | 1 | 4 | 5 | 9 | 2 | 3 | 7 | 10 | 4 | 1 | 7 | 4.81 |
B.2.4 Real application: maximizing the silhouette score
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini G. silhouette | 4.62 | 0.3410 | 0.3080 | 4.60 | 0.6247 | 0.6275 | 2.20 | 0.4610 | 0.5560 | 1.76 | 0.5581 | 0.3132 | 1.88 | 0.6273 | 0.5312 | 2.34 | 0.4198 | 0.3443 | 2.82 | 0.2689 | 0.0295 | 1.10 | 0.6991 | 0.0432 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |
| Iris | Wine | Breast-Cancer | Sonar | QSAR | Vehicle | Australian | Glass | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 3 3.52 | 0.0998 | 0.1 | 3 4.10 | 0.1913 | 0.1738 | 2 2.70 | 0.9234 | 0.9157 | 2 1.22 | 0.6339 | 0.5956 | 3 3.12 | 0.3741 | 0.2978 | 3 4.52 | 0.2946 | 0.2801 | 2 1.66 | 0.7693 | 0.6145 | 6 2.54 | 0.1492 | 0.0699 |
| Gini G. Silhouette | 3 1.40 | 0.08 | 0.1 | 3 1.72 | 0.0902 | 0.0890 | 2 1.12 | 0.8840 | 0.8577 | 2 2.38 | 0.5746 | 0.5423 | 3 1.88 | 0.3485 | 0.2670 | 3 1.10 | 0.2207 | 0.2186 | 2 1.10 | 0.7609 | 0.5797 | 6 1.14 | 0.0798 | 0.0420 |
| Gini | 3 2 | 0.0828 | 0.1 | 3 2 | 0.1172 | 0.1005 | 2 2 | 0.9129 | 0.9052 | 2 2 | 0.5985 | 0.5618 | 3 2 | 0.3419 | 0.2644 | 3 2 | 0.2596 | 0.2517 | 2 2 | 0.6891 | 0.6101 | 6 2 | 0.0584 | 0.0327 |
| Euclidean | 3 - | 0.0757 | 0.0933 | 3 - | 0.0610 | 0.0502 | 2 - | 0.8802 | 0.8524 | 2 - | 0.5636 | 0.5373 | 3 - | 0.3520 | 0.2349 | 3 - | 0.2088 | 0.2269 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0388 | 0.0186 |
| Manhattan | 3 - | 0.0788 | 0.1 | 3 - | 0.0680 | 0.0613 | 2 - | 0.8782 | 0.8489 | 2 - | 0.5818 | 0.5518 | 3 - | 0.3531 | 0.2567 | 3 - | 0.2057 | 0.2234 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0628 | 0.0279 |
| Minkowski | 3 3 | 0.0774 | 0.0933 | 3 3 | 0.0643 | 0.0502 | 2 3 | 0.8812 | 0.8542 | 2 3 | 0.5784 | 0.5419 | 3 3 | 0.3498 | 0.2311 | 3 3 | 0.2100 | 0.2281 | 2 3 | 0.4873 | 0.5590 | 6 3 | 0.0359 | 0.0186 |
| Cosine | 3 - | 0.0188 | 0.02 | 3 - | 0.1370 | 0.1119 | 2 - | 0.9128 | 0.9051 | 2 - | 0.5627 | 0.5325 | 3 - | 0.3945 | 0.3055 | 3 - | 0.1984 | 0.2186 | 2 - | 0.6770 | 0.6725 | 6 - | 0.0599 | 0.0280 |
| Canberra | 3 - | 0.0386 | 0.0467 | 3 - | 0.3990 | 0.3430 | 2 - | 0.9237 | 0.9227 | 2 - | 0.5407 | 0.5225 | 3 - | 0.4449 | 0.3979 | 3 - | 0.2292 | 0.2234 | 2 - | 0.5208 | 0.5203 | 6 - | 0.1123 | 0.0421 |
| Hellinger | 3 - | 0.2906 | 0.2933 | 3 - | 0.2646 | 0.2740 | 2 - | 0.7313 | 0.7156 | 2 - | 0.3424 | 0.4901 | 3 - | 0.5476 | 0.3854 | 3 - | 0.3701 | 0.3275 | 2 - | 0.6701 | 0.5725 | 6 - | 0.1579 | 0.1728 |
| JensenShannon | 3 - | 0.4642 | 0.46 | 3 - | 0.0543 | 0.1352 | 2 - | 0.1272 | 0.3268 | 2 - | 0.5549 | 0.5531 | 3 - | 0.5241 | 0.5378 | 3 - | 0.3695 | 0.4196 | 2 - | 0.6830 | 0.6319 | 6 - | 0.0 | 0.0 |
| PearsonChi2 | 3 - | 0.2688 | 0.2867 | 3 - | 0.4505 | 0.4865 | 2 - | 0.8890 | 0.8683 | 2 - | 0.5973 | 0.5823 | 3 - | 0.3998 | 0.2990 | 3 - | 0.3644 | 0.3509 | 2 - | 0.6705 | 0.5913 | 6 - | 0.2025 | 0.1457 |
| VicisSymmetric1 | 3 - | 0.0667 | 0.0333 | 3 - | 0.2341 | 0.2037 | 2 - | 0.8656 | 0.8420 | 2 - | 0.4488 | 0.5296 | 3 - | 0.4766 | 0.2786 | 3 - | 0.4819 | 0.4031 | 2 - | 0.6502 | 0.5652 | 6 - | 0.1515 | 0.1260 |
| Hassanat | 3 - | 0.0497 | 0.06 | 3 - | 0.3990 | 0.3598 | 2 - | 0.9240 | 0.9157 | 2 - | 0.5834 | 0.5470 | 3 - | 0.3772 | 0.2490 | 3 - | 0.3563 | 0.2931 | 2 - | 0.7858 | 0.7739 | 6 - | 0.0553 | 0.0279 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini G. Silhouette | 8 | 1 | 4 | 2 | 6 | 6 | 9 | 3 | 5 | 8 | 6 | 6 | 11 | 8 | 2 | 5 | 5.62 |
| Gini | 6 | 8 | 3 | 5 | 3 | 7 | 12 | 6 | 4 | 7 | 3 | 1 | 12 | 6 | 3 | 8 | 5.88 |
| Euclidean | 10 | 4 | 5 | 8 | 4 | 11 | 3 | 10 | 8 | 11 | 8 | 7 | 9 | 10 | 10 | 10 | 8 |
| Manhattan | 7 | 3 | 5 | 10 | 2 | 9 | 1 | 10 | 6 | 9 | 9 | 4 | 8 | 11 | 10 | 6 | 6.88 |
| Minkowski | 9 | 2 | 5 | 7 | 5 | 10 | 5 | 10 | 7 | 10 | 7 | 5 | 10 | 9 | 10 | 11 | 7.62 |
| Cosine | 5 | 6 | 10 | 3 | 1 | 8 | 2 | 8 | 12 | 6 | 4 | 8 | 6 | 12 | 5 | 7 | 6.44 |
| Canberra | 2 | 7 | 2 | 1 | 9 | 1 | 11 | 5 | 11 | 2 | 2 | 10 | 4 | 7 | 9 | 4 | 5.44 |
| Hellinger | 1 | 5 | 8 | 9 | 8 | 5 | 8 | 1 | 2 | 4 | 11 | 12 | 1 | 2 | 7 | 2 | 5.38 |
| Jensen-Shannon | 11 | 12 | 9 | 12 | 10 | 12 | 10 | 1 | 1 | 12 | 12 | 9 | 2 | 3 | 4 | 12 | 8.25 |
| PearsonChi2 | 11 | 11 | 12 | 6 | 11 | 3 | 4 | 7 | 3 | 1 | 5 | 2 | 5 | 4 | 6 | 1 | 5.75 |
| VicisSymmetric1 | 3 | 10 | 11 | 11 | 12 | 2 | 7 | 9 | 9 | 5 | 10 | 11 | 3 | 1 | 8 | 3 | 7.19 |
| Hassanat | 4 | 9 | 1 | 4 | 7 | 4 | 6 | 4 | 10 | 2 | 1 | 3 | 7 | 5 | 1 | 9 | 4.81 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini G. Silhouette | 10 | 1 | 2 | 3 | 3 | 8 | 12 | 8 | 4 | 9 | 6 | 6 | 7 | 11 | 6 | 5 | 6.31 |
| Gini | 8 | 8 | 3 | 2 | 2 | 7 | 7 | 6 | 4 | 8 | 3 | 2 | 8 | 6 | 4 | 6 | 5.25 |
| Euclidean | 12 | 4 | 5 | 5 | 4 | 10 | 9 | 10 | 7 | 11 | 8 | 8 | 11 | 8 | 9 | 10 | 8.19 |
| Manhattan | 9 | 2 | 5 | 8 | 1 | 8 | 11 | 10 | 4 | 10 | 9 | 4 | 9 | 9 | 9 | 8 | 7.25 |
| Minkowski | 11 | 2 | 5 | 7 | 6 | 11 | 6 | 10 | 7 | 11 | 7 | 7 | 12 | 7 | 9 | 10 | 8 |
| Cosine | 5 | 5 | 11 | 4 | 5 | 6 | 2 | 4 | 12 | 7 | 4 | 9 | 4 | 11 | 2 | 7 | 6.12 |
| Canberra | 3 | 6 | 4 | 9 | 10 | 3 | 10 | 2 | 10 | 3 | 1 | 11 | 2 | 9 | 12 | 4 | 6.19 |
| Hellinger | 2 | 7 | 12 | 10 | 9 | 4 | 4 | 9 | 2 | 4 | 11 | 12 | 3 | 4 | 7 | 1 | 6.31 |
| Jensen-Shannon | 6 | 9 | 8 | 12 | 12 | 12 | 5 | 7 | 1 | 6 | 12 | 3 | 1 | 1 | 3 | 12 | 6.88 |
| PearsonChi2 | 6 | 12 | 10 | 6 | 11 | 2 | 1 | 3 | 3 | 1 | 5 | 1 | 5 | 3 | 5 | 2 | 4.75 |
| VicisSymmetric1 | 1 | 11 | 9 | 11 | 8 | 5 | 8 | 1 | 11 | 5 | 10 | 10 | 6 | 2 | 8 | 3 | 6.81 |
| Hassanat | 4 | 10 | 1 | 1 | 7 | 1 | 3 | 5 | 9 | 2 | 2 | 5 | 10 | 5 | 1 | 8 | 4.62 |
B.3 Additional experiments on agglomerative hierarchical clustering with 10% level of noise
| Ionosphere | Banknote | German | Heart | Balance | Haberman | Wholesale | Indian Liver Patient | |||||||||||||||||
| Distance | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | #P | #R | ||||||||
| Gini Generalized | 3 1.10 | 0.4114 | 0.6410 | 3 5.07 | 0.8009 | 0.6888 | 3 1.38 | 0.5402 | 0.5380 | 3 2.78 | 0.5563 | 0.5974 | 3 1.35 | 0.6726 | 0.6063 | 3 1.65 | 0.7003 | 0.7353 | 3 3.76 | 0.8622 | 0.8182 | 3 1.10 | 0.5132 | 0.7150 |
| Gini | 3 | 0.4114 | 0.6410 | 3 | 0.3885 | 0.5503 | 3 | 0.5117 | 0.5070 | 3 | 0.4324 | 0.5380 | 3 | 0.5973 | 0.6416 | 3 | 0.5409 | 0.7353 | 3 | 0.4582 | 0.6727 | 3 | 0.5132 | 0.7150 |
| Euclidean | 3 | 0.4114 | 0.6410 | 3 | 0.7538 | 0.6618 | 3 | 0.3190 | 0.4850 | 3 | 0.3424 | 0.5446 | 3 | 0.5566 | 0.5840 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Euclidean Ward | 3 | 0.6712 | 0.6667 | 3 | 0.7137 | 0.6550 | 3 | 0.5420 | 0.5435 | 3 | 0.3074 | 0.5545 | 3 | 0.7044 | 0.6298 | 3 | 0.5699 | 0.7549 | 3 | 0.8195 | 0.7891 | 3 | 0.5880 | 0.7668 |
| Manhattan | 3 | 0.5322 | 0.6439 | 3 | 0.7573 | 0.6720 | 3 | 0.4272 | 0.5110 | 3 | 0.3424 | 0.5446 | 3 | 0.6252 | 0.6336 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Minkowski | 3 3 | 0.4114 | 0.6410 | 3 3 | 0.7385 | 0.6684 | 3 3 | 0.4131 | 0.4920 | 3 3 | 0.3424 | 0.5446 | 3 3 | 0.5247 | 0.5649 | 3 3 | 0.5409 | 0.7353 | 3 3 | 0.4592 | 0.6773 | 3 3 | 0.5132 | 0.7150 |
| Cosine | 3 | 0.4114 | 0.6410 | 3 | 0.6432 | 0.6443 | 3 | 0.4021 | 0.4770 | 3 | 0.3630 | 0.5479 | 3 | 0.6220 | 0.6160 | 3 | 0.5409 | 0.7353 | 3 | 0.5899 | 0.7249 | 3 | 0.5132 | 0.7150 |
| Canberra | 3 | 0.6065 | 0.6496 | 3 | 0.6126 | 0.5685 | 3 | 0.2277 | 0.4770 | 3 | 0.5402 | 0.6502 | 3 | 0.6541 | 0.6335 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Hellinger | 3 | 0.4114 | 0.6410 | 3 | 0.7139 | 0.6553 | 3 | 0.3967 | 0.4770 | 3 | 0.3647 | 0.5479 | 3 | 0.6282 | 0.6385 | 3 | 0.5409 | 0.7353 | 3 | 0.5768 | 0.6818 | 3 | 0.5132 | 0.7150 |
| JensenShannon | 3 | 0.4114 | 0.6410 | 3 | 0.3089 | 0.5554 | 3 | 0.2277 | 0.4770 | 3 | 0.3649 | 0.5479 | 3 | 0.6515 | 0.6000 | 3 | 0.5409 | 0.7353 | 3 | 0.5783 | 0.7158 | 3 | 0.5132 | 0.7150 |
| PearsonChi2 | 3 | 0.6615 | 0.6467 | 3 | 0.5251 | 0.5401 | 3 | 0.3848 | 0.4680 | 3 | 0.3424 | 0.5446 | 3 | 0.5378 | 0.5760 | 3 | 0.6036 | 0.7418 | 3 | 0.5377 | 0.6795 | 3 | 0.5132 | 0.7150 |
| VicisSymmetric1 | 3 | 0.6579 | 0.6581 | 3 | 0.5517 | 0.5816 | 3 | 0.3188 | 0.4780 | 3 | 0.4889 | 0.5446 | 3 | 0.6336 | 0.6096 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Hassanat | 3 | 0.7739 | 0.6496 | 3 | 0.4520 | 0.5904 | 3 | 0.3807 | 0.4820 | 3 | 0.6616 | 0.7096 | 3 | 0.6426 | 0.5920 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 7 | 1 | 2 | 2 | 2 | 1 | 1 | 2 | 1 | 2 | 1 | 1 | 2 | 2 | 1 | 3 | 1.94 |
| Gini | 7 | 12 | 3 | 5 | 10 | 4 | 13 | 2 | 3 | 6 | 3 | 6 | 5 | 3 | 7 | 9 | 6.12 |
| Euclidean | 7 | 3 | 10 | 9 | 11 | 4 | 7 | 2 | 4 | 9 | 6 | 3 | 10 | 5 | 5 | 2 | 6.06 |
| Euclidean Ward | 2 | 6 | 1 | 13 | 1 | 3 | 2 | 1 | 7 | 1 | 2 | 9 | 1 | 1 | 4 | 4 | 3.62 |
| Manhattan | 6 | 2 | 4 | 9 | 8 | 4 | 7 | 2 | 2 | 7 | 6 | 4 | 10 | 11 | 3 | 1 | 5.38 |
| Minkowski | 7 | 4 | 5 | 9 | 13 | 4 | 7 | 2 | 9 | 7 | 6 | 10 | 10 | 7 | 2 | 5 | 6.69 |
| Cosine | 7 | 7 | 6 | 8 | 9 | 4 | 3 | 2 | 10 | 11 | 5 | 10 | 9 | 8 | 8 | 8 | 7.19 |
| Canberra | 5 | 8 | 12 | 3 | 3 | 4 | 7 | 2 | 11 | 10 | 11 | 2 | 13 | 12 | 8 | 6 | 7.31 |
| Hellinger | 7 | 5 | 7 | 7 | 7 | 4 | 5 | 2 | 5 | 5 | 9 | 10 | 3 | 9 | 6 | 10 | 6.31 |
| JensenShannon | 7 | 13 | 12 | 6 | 4 | 4 | 4 | 2 | 11 | 13 | 11 | 10 | 7 | 13 | 8 | 13 | 8.62 |
| PearsonChi2 | 3 | 10 | 8 | 9 | 12 | 2 | 6 | 2 | 5 | 3 | 10 | 7 | 3 | 9 | 8 | 11 | 6.75 |
| VicisSymmetric1 | 4 | 9 | 11 | 4 | 6 | 4 | 7 | 2 | 11 | 12 | 11 | 5 | 8 | 6 | 8 | 12 | 7.5 |
| Hassanat | 1 | 11 | 9 | 1 | 5 | 4 | 7 | 2 | 8 | 4 | 4 | 8 | 6 | 4 | 8 | 7 | 5.56 |
| Distances | I | B | G | He | Ba | Ha | W | In | Ir | Wi | Bc | S | QS | V | A | Gl | Rank |
| Gini Generalized | 7 | 1 | 2 | 3 | 8 | 3 | 1 | 2 | 1 | 3 | 1 | 3 | 4 | 9 | 3 | 1 | 3.25 |
| Gini | 7 | 12 | 4 | 13 | 1 | 3 | 13 | 2 | 3 | 5 | 4 | 9 | 5 | 4 | 13 | 10 | 6.75 |
| Euclidean | 7 | 4 | 6 | 8 | 11 | 3 | 7 | 2 | 5 | 10 | 6 | 9 | 10 | 11 | 3 | 5 | 6.69 |
| Euclidean Ward | 1 | 6 | 1 | 4 | 5 | 1 | 2 | 1 | 2 | 1 | 2 | 1 | 1 | 6 | 6 | 2 | 2.62 |
| Manhattan | 6 | 2 | 3 | 8 | 3 | 3 | 7 | 2 | 4 | 8 | 6 | 11 | 10 | 11 | 5 | 4 | 5.81 |
| Minkowski | 7 | 3 | 5 | 8 | 13 | 3 | 7 | 2 | 10 | 8 | 6 | 3 | 10 | 5 | 2 | 3 | 5.94 |
| Cosine | 7 | 7 | 9 | 5 | 6 | 3 | 3 | 2 | 7 | 11 | 5 | 3 | 9 | 7 | 7 | 9 | 6.25 |
| Canberra | 3 | 10 | 9 | 2 | 4 | 3 | 7 | 2 | 11 | 6 | 11 | 2 | 13 | 10 | 7 | 12 | 7 |
| Hellinger | 7 | 5 | 9 | 5 | 2 | 3 | 5 | 2 | 7 | 7 | 9 | 3 | 7 | 2 | 1 | 7 | 5.06 |
| JensenShannon | 7 | 11 | 9 | 5 | 9 | 3 | 4 | 2 | 11 | 12 | 11 | 3 | 2 | 13 | 7 | 8 | 7.31 |
| PearsonChi2 | 5 | 13 | 13 | 8 | 12 | 2 | 6 | 2 | 7 | 4 | 10 | 13 | 7 | 2 | 7 | 6 | 7.31 |
| VicisSymmetric1 | 2 | 9 | 8 | 8 | 7 | 3 | 7 | 2 | 11 | 13 | 11 | 11 | 3 | 8 | 7 | 11 | 7.56 |
| Hassanat | 3 | 8 | 7 | 1 | 10 | 3 | 7 | 2 | 6 | 2 | 3 | 8 | 6 | 1 | 7 | 13 | 5.44 |